### v1

# When a Reinforcement Learning Agent Encounters Unknown Unknowns

Abstract
An AI agent might surprisingly find she has reached an unknown state which she has never been aware of – an unknown unknown. We mathematically ground this scenario in reinforcement learning: an agent, after taking an action calculated from value functions $Q$ and $V$ defined on the aware domain, reaches a state out of the domain. To enable the agent to handle this scenario, we propose an episodic Markov decision process with growing awareness (EMDP-GA) model, taking a new noninformative value expansion (NIVE) approach to expand value functions to newly aware areas: when an agent arrives at an unknown unknown, value functions $Q$ and $V$ whereon are initialised by noninformative beliefs – the averaged values on the aware domain. This design is out of respect for the complete absence of knowledge in the newly discovered state. The upper confidence bound momentum Q-learning is then adapted to the growing awareness for training the EMDP-GA model. We prove that (1) the regret of our approach is asymptotically consistent with the state of the art (SOTA) without exposure to unknown unknowns in an extremely uncertain environment, and (2) our computational complexity and space complexity are comparable with the SOTA – these collectively suggest that though an unknown unknown is surprising, it will be asymptotically properly discovered with decent speed and an affordable cost.
1 Introduction
[…]
Consider a scenario that is completely untouched in the literature of reinforcement learning – an agent might surprisingly find she has reached an unknown state where she is even unaware of. The former US Secretary of Defense Donald Rumsfeld referred to this scenario as an unknown unknown in a news conference press . Unknown unknowns broadly exist in security-critical applications, such as autonomous vehicles van2018autonomous , quantitative finance liu2021finrl and economics chang2020decentralized , when the environment is extremely under-explored and uncertain.
Figure 1: Growing awareness
This paper proposes the first mathematical model to ground unknown unknowns in reinforcement learning. Although actions are calculated by maximising value functions $Q$ and $V$ that are defined on the aware domain, some of them could move the agent to unknown unknown states in the unaware domain. Based upon this, this paper designs a novel model, episodic Markov decision process with growing awareness (EMDP-GA), which employs a newly designed noninformative value expansion (NIVE) approach to expand value functions to the newly aware states, when the agent surprisingly reaches there. Figure 1 illustrates the expansion of the aware domain, where $\mathcal{S}$ is the whole state space and $\mathcal{S}_{t}$ is the aware domain after exploring $t$ episodes. The initial values are chosen as a noninformative prior tibshirani1989noninformative , such as the averaged values on the previous aware domain, out of respect for the extreme lack of knowledge in this unknown unknown state. We further adapt upper confidence bound momentum Q-learning (UCBMQ) menard2021ucb to the growing awareness for training the EMDP-GA, termed as UCBMQ-GA. UCBMQ employs a momentum term to correct the bias in Q-learning, but relies on the upper bounds of value functions $Q$ and $V$. We prove that our NIVE can preserve this ‘upper-bound’ property when expanding the value functions, which will be invalidated by multiplying NIVE’s expansion by any constant $d<1$. This confirms the efficiency of NIVE as well as our adaptation made in EMDP-GA and UCBMQ-GA.
Extensive theoretical analysis is conducted for verifying our approach. We analyse the regret over the whole course, which is defined as the cumulative reward loss over the whole learning process against the hidden optimal policy. The regret has an upper bound in order of $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA+H^{2}S\sqrt{T})$, where $S$ is the size of the state space, $A$ is the size of the action space, $T$ is the number of episodes, and $H$ is the length of an episode. This regret bound is sublinear with respect to $T$, and comparable with the state of the art (SOTA), including upper confidence bound-advantage zhang2020almost , upper confidence bound momentum Q-learning menard2021ucb and monotonic value propagation zhang2024settling , despite our exposure to unknown unknowns, while they are not. The computing complexity and space complexity of employing the UCBMQ-GA approach for training the EMDP-GA model are of order $\mathcal{O}(H(S+A)T)$ and $\mathcal{O}(HS^{2}A)$, respectively, which are comparable with the SOTA menard2021ucb . These results collectively suggest that although unknown unknowns are surprising, they can nonetheless be discovered asymptotically properly with a decent speed and an affordable cost.
2 Related Works
Episodic Markov decision process (EMDP)
A large volume of literature has been on developing EMDP over the last decade. Neu et al. study the adversarial stochastic shortest path problem in EMDPs neu2012adversarial . Zimin and Neu investigate EMDPs with a layered structure zimin2013online . Dann and Brunskill study EMDPs from the perspective of probably approximately correct (PAC) learning dann2015sample . EMDP algorithms can be roughly classified into model-based and model-free approaches. Model-based approaches learn the EMDP model prior to or concurrently with policy optimisation and require $\Omega(S^{2}AH)$ memory space to store the value function estimators moerland2023model , in which recent developments have achieved lower regret and burn-in costs azar2017minimax ; dann2019policy ; zanette2019tighter ; efroni2019tight ; zhang2020almost ; zhang2021reinforcement ; zhang2024settling ; li2024breaking . Here, burn-in costs are the minimum sample size needed for an algorithm to operate sample-optimally. In contrast, model-free approaches estimate the optimal value function or optimise the policy directly, without explicit model estimation, and require only $O(SAH)$ memory to store these estimators. Extensive theoretical analyses of model-free approaches have also been conducted zhang2020almost ; menard2021ucb ; mao2021near ; dann2021provably ; tiapkin2022dirichlet ; li2023breaking ; agrawal2024optimistic . Amongst them, a remarkable result is a minimax regret lower bound, $\mathop{\min}\{\sqrt{SAH^{3}T},HT\}$, established by Domingues et al. domingues2021episodic , offering a theoretical benchmark.
Awareness and reverse Bayes
Similar terms are seen in the literature of choice theory grant2015preference ; piermont2017introspective ; kochov2018behavioral . Reverse Bayesianism is proposed to model the expanding decision universe as awareness grows karni2013reverse ; karni2015probabilistic ; karni2017awareness . Belief formation and responses to unforeseen events, as predicted by reverse Bayesianism, are experimentally studied becker2020reverse . This framework was then applied in economic theory viero2021intertemporal ; chakravarty2022reverse ; karni2021reverse and epistemology zednik2016bayesian . However, no practical approach has been proposed to train an agent to be able to handle unknown unknowns.
3 Notations and Preliminaries
Episodic Markov decision process (EMDP))
An EMDP is defined by a tuple $M=(\mathcal{S},\mathcal{A},P,r,H,T)$ fiechter1994efficient ; kaufmann2021adaptive , described below. $\mathcal{S}$ and $\mathcal{A}$ are finite state space and finite action space, respectively; mathematically, they are subsets of Euclidean spaces. If an agent at a state $s\in\mathcal{S}$ takes an action $a\in\mathcal{A}$, it will be transited to a new state $s^{\prime}\in\mathcal{S}$. An EMDP agent takes episodic actions for learning – after an episode, the agent comes back to the initial state $s_{1}$, and starts the next episode. Usually, episodes involve the same number of time steps, termed as horizon $H\in\mathbb{N}$. $T\in\mathbb{N}$ is the number of episodes in a learning process. We can thus denote the episode $t$ by the trajectory $(s_{1}^{t},a_{1}^{t},\cdots,s_{H}^{t},a_{H}^{t},s_{H+1}^{t})$. For any episode, at time step $h\in[H]\triangleq\{1,\cdots,H\}$, a transition function $P_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ defines the probability $P_{h}(s^{\prime}|s,a)$ of transition from state $s$ to state $s^{\prime}$ when taking action $a$. This interaction leads to an immediate reward $r_{h}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. Let $\pi_{h}(a|s)$ denote the probability of taking action $a$ at state $s$. A policy model $\pi=(\pi_{1},\cdots,\pi_{H})\in\Pi$ characterises the agent’s action course. All possible policy models constitute a policy space $\Pi=\left\{\pi=(\pi_{1},\cdots,\pi_{H})|\pi_{h}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]\right\}$. We denote the policy used to get episode $t$ by $\pi^{t}\in\Pi$.
Value functions and regret
At time step $h\in[H]$, for any state $s\in\mathcal{S}$ and action $a\in\mathcal{A}$, the Q-value function of a policy $\pi$ is defined as $Q_{h}^{\pi}(s,a)=r_{h}(s,a)+\mathbb{E}_{p,\pi}\left[\sum_{l=h+1}^{H}r_{l}(s_{l},a_{l})|s_{h}=s,a_{h}=a\right]$, and the value function is defined as $V_{h}^{\pi}(s)=\mathbb{E}_{p,\pi}\left[\sum_{l=h}^{H}r_{l}(s_{l},a_{l})|s_{h}=s\right]$. Their relationship is governed by the Bellman equation
$V_{h}^{\pi}(s)=\sum_{a\in A}\pi(a|s)Q_{h}^{\pi}(s,a),\ \ Q_{h}^{\pi}(s,a)=r_{h}(s,a)+\sum_{s^{\prime}\in S}p(s^{\prime}|s,a)V_{h}^{\pi}(s^{\prime}).$
The objective of reinforcement learning is to learn a policy that maximises the cumulative rewards, denoted by $V_{1}^{\pi}(s)$. Shreve and Bertsekas prove that, under certain conditions, there exists an optimal deterministic policy $\pi$ satisfy $V_{h}^{*}(s)\triangleq V_{h}^{\pi^{*}}(s)=\mathop{\max}_{\pi}V_{h}^{\pi}(s),\ Q_{h}^{*}(s,a)\triangleq Q_{h}^{\pi^{*}}(s,a)=\mathop{\max}_{\pi}Q_{h}^{\pi}(s,a)$ shreve1978alternative . This result leads to the Bellman optimality equation:
$V_{h}^{*}(s)=\mathop{\max}_{a\in A}Q_{h}^{*}(s,a),\ \ Q_{h}^{*}(s,a)=r_{h}(s,a)+\sum_{s^{\prime}\in S}p(s^{\prime}|s,a)V_{h}^{*}(s^{\prime}).$
The performance of an algorithm is often evaluated by its regret against the optimal policy $\mathcal{R}^{T}=\sum_{t=1}^{T}\left[V^{*}_{1}(s_{1})-V_{1}^{\pi^{t}}(s_{1})\right]$. Note that $V_{1}^{\pi^{t}}(s_{1})\leq V^{*}_{1}(s_{1})$ for any $t\in[T]$.
Upper confidence bound momentum Q-learning (UCBMQ)
UCBMQ is designed for training an EMDP model menard2021ucb . Its objective is to approximate the optimal policy by controlling the regret. It optimises four functions: estimated Q-value function $Q_{h}^{t}(s,a)$, bias-value function $V_{h,s,a}^{t}(s^{\prime})$, upper bound on the optimal Q-value function $\overline{Q}_{h}^{t}(s,a)$, and upper bound on the optimal value function $\overline{V}_{h}^{t}(s)$. In iteration $t$, the agent interacts with the environment to collect episode $t$, updating the visitation counts $n_{h}^{t}(s,a)$ for each state-action pair $(s,a)$. Next, we update the count-dependent parameters, including the bonus term $\beta_{h}^{t}(s,a)$. Finally, we update the four value functions. The core method of updating $Q_{h}^{t}(s,a)$ retains the standard Q-learning update, which is inspired by the Bellman optimality equation, and adds a momentum correction to remove the bias. $V_{h,s,a}^{t}(s^{\prime})$ is updated as a convex combination of the previous bias-value and the optimistic next-step value, $\overline{Q}_{h}^{t}(s,a)$ is updated by adding $\beta_{h}^{t}(s,a)$ to $Q_{h}^{t}(s,a)$ and $\overline{V}_{h}^{t}(s)$ is updated according to Bellman optimality equation. The procedure of UCBMQ is described in Appendix B. UCBMQ can guarantee a regret bound of $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA)$, where $H$, $S$, $A$ and $T$ are the horizon, the size of the state space, the size of the action space and the number of learning episodes, respectively. This regret bound matches the lower bound $\Omega(\sqrt{SAH^{3}T})$ domingues2021episodic for large enough $T$.
4 Growing Awareness in Extremely Under-explored, Uncertain Environment
The awareness of an agent can be considerably restricted and problematic in an extremely under-explored and uncertain environment. We are aiming to grow the awareness along the agent’s course of exploring the environment.
4.1 Mathematically Grounding Awareness in Reinforcement Learning
An example of a spacecraft
Consider a spacecraft controlled by a reinforcement learning agent. She has explored her homeland (or home planet), which is not necessarily fully completed. The agent’s value functions are naturally defined on her homeland, which is her current aware domain. After taking an action calculated from the value functions defined on the aware domain, the spacecraft may go out of her home planet. The spacecraft then starts to explore the previously unaware domain. The exploration requires an initialisation for the value functions to kick off. This calls for an expansion strategy for the value functions to newly discovered states. The value functions are optimised afterwards. After sufficient exploration, a good spacecraft is expected to develop an (near-)optimal policy – this is exactly the goal of this paper. Figure 2 illustrates this example.
Mathematically, we translate awareness to the domain of value functions, as the definitions below.
Definition 1 (aware domain).
The value functions are defined over the awareness domain. In this domain, the agent is aware of each state’s existence but needs to fully explore it to ascertain its value.
Definition 2 (unaware domain).
The unaware domain comprises the states whose existence and values are both unknown to the agent.
Remark 1.
In our framework, once a state in the unaware domain is explored, it moves to the aware domain and never reverts to the unaware domain. States in the awareness domain remain there permanently. Hence, the agent’s awareness monotonically increases without forgetting previously discovered states.
4.2 Episodic Markov Decision Processes with Growing Awareness
We now define our episodic Markov decision processes with growing awareness (EMDP-GA) model as a tuple $M_{GA}=(\mathcal{S},\mathcal{S}_{0},\mathcal{A},P,r,H,T)$, where $\mathcal{S}_{0}\subseteq\mathcal{S}$ is the initial aware domain. Note that the initial state of each learning episode is fixed as $s_{1}\in\mathcal{S}_{0}$. If the agent reaches an unknown unknown state $s$ after taking an action, she will be aware of the existence of $s$, so that the aware domain grows by the state $s$. Let $\mathcal{S}_{t}$ denote the aware domain after exploring $t$ episodes. At episode $t+1$, the agent’s trajectory is $\left(s^{t+1}_{1},\cdots,s^{t+1}_{H}\right)$, and the aware domain is updated accordingly as $\mathcal{S}_{t+1}=\mathcal{S}_{t}\cup\left\{s_{1}^{t+1},\cdots,s_{H}^{t+1}\right\}$. Note that $s_{1}\in\mathcal{S}_{0}\subseteq\cdots\subseteq\mathcal{S}_{T}\subseteq\mathcal{S}$.
Figure 2: Illustration of a spacecraft
After $t$ episodes, we denote by $U_{t}$ the unknown unknown states. The extended aware domain is thus $\tilde{\mathcal{S}}_{t}=\mathcal{S}_{t}\cup U_{t}$. EMDP usually defines state space and action space independently. Inheriting this convention, in this paper, the agent selects actions from a fixed action space $\mathcal{A}$, despite the fact that the aware domain is growing, even when the agent is placed at new states. Note that the policy $\pi^{t}$, which is used to get episode $t$, is defined on the aware domain after exploring $t-1$ episodes. Thus, the policy $\pi^{t}=(\pi_{1}^{t},\cdots,\pi_{H}^{t})$ is defined on $\tilde{\mathcal{S}}_{t-1}$; i.e., $\pi_{h}^{t}:\tilde{\mathcal{S}}_{t-1}\times\mathcal{A}\rightarrow[0,1]$ for each $h$. Here, $\pi_{h}^{t}(\cdot|s)$ is a distribution over $\mathcal{A}$. Consequently, $\pi^{t}$ effectively constitutes a policy over the entire state space $\mathcal{S}$.
Homeland condition
Exploring in an extremely unexplored, uncertain world requires a ‘homeland’. Space exploration needs an earth. We assume that with high probability, unknown unknowns are no more valuable than the average on known states; a very high-value target still exists but with a high risk. Mathematically, with probability at least $1-\frac{\delta}{3}$, the average value of the known states exceeds that of any unknown unknown state:
$\forall s\in\mathcal{S}-\mathcal{S}_{t},Q^{*}_{h}(s,a)\leq\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}Q^{*}_{h}(\hat{s},a),\ V^{*}_{h}(s)\leq\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}V^{*}_{h}(\hat{s}).$
It coincides with the fact that exploring in an extremely under-explored, uncertain environment is a high-risk, high-gain mission.
4.3 Noninformative Prior for Domain Expansion to Newly Aware States
When an agent encounters an unknown unknown state that lies outside the currently aware domain, the reinforcement learning algorithm becomes invalid. The unknown unknown status implies that it exhibits properties that significantly differ from any known state, and prior information is completely absent. Out of respect for this extreme lack of knowledge, we design a noninformative value expansion (NIVE) approach to expand value functions to the newly discovered states – the value functions are initialised using the averages over the current aware domain. We introduce NIVE with more details below.
Algorithm of NIVE
NIVE updates three value functions: estimated Q-value function $Q_{h}^{t}(s,a)$, bias-value function $V_{h,s,a}^{t}(s^{\prime})$, and upper bound on the optimal value function $\overline{V}_{h}^{t}(s)$. When the aware domain grows from $\mathcal{S}_{t-1}$ to $\mathcal{S}_{t}$, we would like to expand the domains of three value functions: $Q^{t-1}_{h}(s,a)$, $\overline{V}^{t-1}_{h}(s)$ and $V^{t-1}_{h,s,a}(s^{\prime})$, accordingly. NIVE calculates the average values over the aware domain $\mathcal{S}_{t-1}$ as estimators for states in $\mathcal{S}_{t}-\mathcal{S}_{t-1}$:
$Q_{h,avg}^{t}(a)=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}Q_{h}^{t}(s,a),\ \ \ \overline{V}_{h,avg}^{t}=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}\overline{V}_{h}^{t}(s),\ \ \ V_{h,s,a,avg}^{t}=\frac{1}{|\mathcal{S}_{t}|}\sum_{s^{\prime}\in\mathcal{S}_{t}}V_{h,s,a}^{t}(s^{\prime}),$
$V_{h,a,avg}^{t}(s^{\prime})=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}V_{h,s,a}^{t}(s^{\prime}),\ \ \ \text{and}\ \ \ V_{h,a,avg}^{t}=\frac{1}{|\mathcal{S}_{t}|^{2}}\sum_{s,s^{\prime}\in\mathcal{S}_{t}}V_{h,s,a}^{t}(s^{\prime}).$
In many reinforcement learning algorithms, such as UCBMQ menard2021ucb used later, upper bounds $\overline{Q}_{h}^{t}(s,a)$ and $\overline{V}_{h}^{t}(s)$ are used to calculate actions. After the surprising encounter, the NIVE expands the estimators of value functions to newly discovered states by assigning them the average values computed over known states. Crucially, the expanded functions $\tilde{\overline{Q}}_{h}^{t}(s,a)$ and $\tilde{\overline{V}}_{h}^{t}(s)$ remain valid upper bounds for $Q^{*}_{h}(s,a)$ and $V_{h}^{*}(s)$, respectively, as the following lemma. The proof is given in Appendix I (implied in Lemma 9).
Lemma 1.
For any time step $t\in[T]$ in episode $h\in[H]$, the following inequality holds with probability at least $1-\frac{\delta}{2}$,
$\tilde{\overline{Q}}_{h}^{t}(s,a)\geq Q^{*}_{h}(s,a),\ \tilde{\overline{V}}_{h}^{t}(s)\geq V_{h}^{*}(s).$
Figure 3: Illustration of updating functions in UCBMQ-GA
Further, we may prove the following corollary. The proof is given in Appendix E.
Corollary 1.
If one were to use any scalar multiple $d<1$ of the average values to expand the functions, the upper-bound property could become invalid. In some cases, assigning $d\overline{V}_{h,avg}^{t},\ (d\neq 1)$ to a newly discovered state would produce an expanded function that no longer upper bounds $V_{h}^{*}(s)$.
This corollary confirms the efficiency of NIVE – multiplying any constant $d<1$ to NIVE will make it invalid.
4.4 Learning an Agent with Growing Awareness to Unknown Unknowns
We now design an optimisation method for training our EMDP-GA model for obtaining an agent with growing awareness to handle the unknown unknowns. We adapt upper confidence bound momentum Q-learning (UCBMQ) menard2021ucb to the growing awareness, termed as UCBMQ-GA.
Algorithm of UCBMQ-GA
UCBMQ-GA optimises four functions: estimated Q-value function $Q_{h}^{t}(s,a)$, bias-value function $V_{h,s,a}^{t}(s^{\prime})$, upper bound on the optimal Q-value function $\overline{Q}_{h}^{t}(s,a)$, and upper bound on the optimal value function $\overline{V}_{h}^{t}(s)$.
Let $\chi_{h}^{t}(s,a)=\mathbb{I}(s_{h}^{t}=s,a_{h}^{t}=a)$ denote the indicator function for the occurrence of $(s,a)$ at time step $h$ in episode $t$. Thus, $n_{h}^{t}(s,a)=\sum_{l=1}^{t}\chi_{h}^{l}(s,a)$ as the visitation count for the pair $(s,a)$ appearing at time step $h$ over $t$ episodes. Define $\chi_{h}^{t}(s,a)=n_{h}^{t}(s,a)=0$ if state $s\notin\mathcal{S}_{t}$. Following the convention, and to the brevity, we define $0\times\infty=0$ and $\frac{1}{0}=\infty$.
For any transition function $p$ and any function $f:S\rightarrow\mathbb{R}$, we define $pf(s,a)=\mathbb{E}_{s^{\prime}\sim p(\cdot|s,a)}[f(s^{\prime})].$ We denote by $p_{h}^{t}$ the Dirac distribution concentrated at $(s_{h+1}^{t})$, and then, $p_{h}^{t}f(s,a)=f(s_{h+1}^{t})$. The procedure of UCBMQ-GA for training EMDA-GA is also explained below. More details are shown in Appendix C.
(0) Initialisation
For state $s\in\mathcal{S}_{0}\subseteq\mathcal{S}$, we set $V^{0}_{h,s,a}=\overline{V}^{0}_{h}=H$, and $Q^{0}_{h}=0$. Then, go to (1).
(1) Decision making
In episode $t$, if the current state is aware (i.e., $s\in\mathcal{S}_{t-1}$), the agent takes an action that maximises $\overline{Q}_{h}^{t-1}(s,a)$. Conversely, if the agent encounters an unknown unknown state (i.e., $s\notin\mathcal{S}_{t-1}$), it chooses an action that maximises ${Q}_{h,avg}^{t-1}(a)$.
The algorithm terminates here if the termination conditions have been met. Otherwise, if the agent go beyond the aware domain, go to (2); otherwise, go to (3).
(2) Embedded NIVE for growing awareness
After completing an episode, the awareness set is updated from $\mathcal{S}_{t-1}$ to $\mathcal{S}_{t}$. UCBMQ-GA then employs NIVE to expand $\tilde{Q}^{t-1}_{h}(s,a)$, $\tilde{\overline{V}}^{t-1}_{h}(s)$ and $\tilde{V}^{t-1}_{h,s,a}(s^{\prime})$ to the newly discovered area, as follows,
$\tilde{Q}^{t-1}_{h}(s,a)=\begin{cases}Q^{t-1}_{h}(s,a),&s\in\mathcal{S}_{t-1},\\ Q^{t-1}_{h,avg}(a),&s\in\mathcal{S}_{t}-\mathcal{S}_{t-1},\end{cases}$ $\ \ \ \tilde{\overline{V}}^{t-1}_{h}(s)=\begin{cases}\overline{V}^{t-1}_{h}(s),&s\in\mathcal{S}_{t-1},\\ \overline{V}_{h,avg}^{t-1},&s\in\mathcal{S}_{t}-\mathcal{S}_{t-1},\end{cases}\text{ and}$
$\tilde{V}^{t-1}_{h,s,a}(s^{\prime})=\begin{cases}V^{t-1}_{h,s,a}(s^{\prime}),&s,s^{\prime}\in\mathcal{S}_{t-1},\\ V_{h,s,a,avg}^{t-1},&s\in\mathcal{S}_{t-1},s^{\prime}\in\mathcal{S}_{t}-\mathcal{S}_{t-1},\\ V_{h,a,avg}^{t-1}(s^{\prime}),&s\in\mathcal{S}_{t}-\mathcal{S}_{t-1},s^{\prime}\in\mathcal{S}_{t-1},\\ V_{h,a,avg}^{t-1},&s,s^{\prime}\in\mathcal{S}_{t}-\mathcal{S}_{t-1}.\end{cases}$
The exploration bonus is then calculated with the expanded functions: if $n_{h}^{t}(s,a)=0$ or $s\notin\mathcal{S}_{t}$, $\beta_{h}^{t}(s,a)=H$; otherwise,
$\beta_{h}^{t}(s,a)=2\sqrt{\frac{\zeta W_{h}^{t}(s,a)}{n_{h}^{t}(s,a)}}+53H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}+\sum_{k=1}^{t}\frac{\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)}{Hlog(T)n_{h}^{t}(s,a)}p_{h}^{k}\left(\tilde{V}_{h,s,a}^{k-1}-\tilde{\overline{V}}_{h+1}^{k-1}\right)(s,a),$
where $W_{h}^{t}(s,a)=\sum_{k=1}^{t}\frac{\chi_{h}^{k}(s,a)}{n_{h}^{k}(s,a)}p_{h}^{k}\left(\tilde{\overline{V}}_{h+1}^{k-1}-\sum_{l=1}^{t}\frac{\chi_{h}^{l}(s,a)}{n_{h}^{l}(s,a)}p_{h}^{l}\tilde{\overline{V}}_{h+1}^{l-1}\right)^{2}(s,a)$.
Then, go to (3).
(3) Optimising value functions
The value functions $Q_{h}^{t}(s,a)$, $V_{h,s,a}^{t}(s^{\prime})$, $\overline{Q}_{h}^{t}(s,a)$ and $\overline{V}_{h}^{t}(s)$ are then optimised as illustrated in Figure 3. The learning rate and momentum term are chosen as
$\alpha_{h}^{t}(s,a)=\frac{\chi_{h}^{t}(s,a)}{n_{h}^{t}(s,a)},\ \gamma_{h}^{t}(s,a)=\chi_{h}^{t}(s,a)\frac{H}{H+n_{h}^{t}(s,a)}\frac{n_{h}^{t}(s,a)-)}{n_{h}^{t}(s,a)},\text{ and}$
$\eta_{h}^{t}(s,a)=\alpha_{h}^{t}(s,a)+\gamma_{h}^{t}(s,a).$
$\zeta$ is the exploration threshold which helps in regulating the exploration bonus.
Then, go to (1).
5 Theoretical Analysis
This section presents theoretical analysis of our algorithm.
5.1 Asymptotical Consistency of Learning Unknown Unknowns
Oracle agent
The theoretical analysis aims to compare our approach against the (unknown) optimal policy $\pi^{*}=(\pi_{1}^{*},\cdots,\pi_{H}^{*})$, $\pi_{h}^{*}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$. This is characterised by the regret of UCBMQ-GA against this optimal policy, defined as $\mathcal{R}^{T}=\sum_{t=1}^{T}[V^{\pi^{*}}(s_{1})-V^{\pi^{t}}(s_{1})]$. To characterise this optimal policy, we assume the existence of an oracle agent that is aware of the entire state space $\mathcal{S}$ and can maximise the value function accordingly.
We make the following assumption about the transition function.
Assumption 1 (non-zero probability of transition).
For the transition function, We assume that $1-\left(\frac{\delta}{3}\right)^{\frac{1}{H[\sqrt{T}]}}\leq\frac{1}{S}$, and the following holds
$\forall(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S},P(s^{\prime}|s,a)\geq 1-\left(\frac{\delta}{3}\right)^{\frac{1}{H[\sqrt{T}]}}.$
Remark 2.
When $T$ is large enough, $1-\left(\frac{\delta}{3}\right)^{\frac{1}{H[\sqrt{T}]}}\leq\frac{1}{S}$ holds and guarantees that $P(\cdot|s,a)$ can be a distribution over $\mathcal{S}$. Assumption 1 means that the agent can probably reach any state from the current state when taking any action.
We then define an aware moment of a state as follows.
Definition 3 (aware moment).
For any state $s\in\mathcal{S}$, define its aware moment as $t(s)=\mathop{min}\limits_{t\in[T]}\{t:s\in\mathcal{S}_{t}\}$; i.e., the first episode after which the agent is aware of $s$.
Naturally, $t(s)\geq 0$. We then have an upper bound on the aware moment. The proof is detailed in Appendix F.
Theorem 1.
With probability at least $1-\frac{\delta}{3}$, the aware moment is upper-bounded: $t(s)\leq\sqrt{T}$, where $T$ is the number of learning episodes.
Remark 3.
This theorem implies that, with high probability, the agent becomes aware of all states within $\sqrt{T}$ episodes.
This finding is instrumental in deriving regret upper bounds. Then, we obtain the regret bound as follows. A detailed proof is given in Appendix J.
Theorem 2.
We assume that $T>3$ and $\zeta=log\left(\frac{96eHSA(2T+1)}{\delta}\right)$. When the UCNMQ-GA is used to train EMDP-GA, with probability at least $1-\delta$, the regret satisfies
$\mathcal{R}^{T}\leq\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA+H^{2}S\sqrt{T}),$
where $H$ is the horizon, $S$ is the size of the state space, $A$ is the size of the action space and $T$ is the number of learning episodes.
Remark 4.
This regret bound is sublinear with respect to the number of episodes $T$. It matches the SOTA, despite that they are not exposed to unknown unknowns, but we are.
Remark 5.
We have now proved that unknown unknowns could be surprising, but can be asymptotically consistently discovered.
Comparison with existing methods, but without exposure to unknown unknowns.
Compared with SOTA, for $H\leq\frac{A}{S}$, the regret bound of our UCBMQ-GA matches UCBMQ ($\mathcal{O}(\sqrt{SAH^{3}T}+SAH^{4})$) menard2021ucb . For $T\geq H^{5}SA$ and $H\leq\frac{A}{S}$, the regret bound of UCBMQ-GA matches monotonic value propagation (MVP, $\mathcal{O}(\sqrt{H^{3}SAT})$) zhang2024settling . Other methods, upper Bayes-confidence bound value iteration (Bayes-UCBVI) tiapkin2022dirichlet and upper confidence bound-advantage(UCB-Adv) zhang2020almost , are worse than the three methods in either case. Table 1 shows the comparisons.
Table 1: Comparisons with the State of the Art
Algorithm Regret Bound Comput. Complex. Spac. Complex.
UCBMQ menard2021ucb $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA)$ $\mathcal{O}(H(S+A)T)$ $\mathcal{O}(HS^{2}A)$
MVP zhang2024settling $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT})$ $\mathcal{O}(HSAT)$ $\mathcal{O}(HS^{2}A)$
UCB-Adv zhang2020almost $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{\frac{33}{4}}S^{2}A^{\frac{3}{2}}T^{\frac{1}{4}})$ $\mathcal{O}(HAT)$ $\mathcal{O}(HSA)$
Bayes-UCBVI tiapkin2022dirichlet $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{3}S^{2}A)$ $\mathcal{O}(BHS^{2}AT)$ $\mathcal{O}(HS^{2}A)$
UCBMQ-GA (ours) $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA+H^{2}S\sqrt{T})$ $\mathcal{O}(H(S+A)T)$ $\mathcal{O}(HS^{2}A)$
Proof sketch
The proof is in five steps. $\mathring{\overline{Q}}_{h}^{t}(s,a)$ and $\mathring{\overline{V}}_{h}^{t}(s)$ are defined in Appendix A.
Step 1: Upper bound $(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)$
Combining the estimation of $\mathring{\overline{Q}}_{h}^{t}(s,a)$ in Lemma 10 and the upper-bound on $\beta_{h}^{t}(s,a)$ in Lemma 11, we obtain the upper-bound of $(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)$.
Step 2: Upper bound the local optimistic regret $\hat{R}_{h}^{T}(s,a)$
We define the local optimistic regret as
$\hat{R}_{h}^{T}(s,a)\triangleq\sum_{t=0}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)=\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a).$
With the result in Step 1, we can decompose $\hat{R}_{h}^{T}(s,a)$ and upper bound each term.
Step 3: Replace $\chi_{h}^{t}$ with $\overline{p}_{h}^{t}$ in the upper-bound on $\hat{R}_{h}^{T}(s,a)$
Following Step 2, we modify the upper bound on the local optimistic regret using $\overline{p}_{h}^{t}(s,a)$, the probability to reach $(s,a)$ at time step $h$ in the episode $t$ as below,
$\hat{R}_{h}^{T}(s,a)\leq$ $63log(T)\sqrt{\zeta\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}+1754H^{3}log(T)^{2}\zeta$
$+(1+\frac{83}{H})\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a).$
Step 4: Upper bound the regret $\hat{R}_{h}^{T}$ at step $h$
We define the step $h$ regret as $\hat{R}_{h}^{T}\triangleq\sum_{s\in S}\sum_{t=0}^{T-1}\overline{p}_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)$. For each $s$, the composition of the step $h$ regret follows by $t(s)$. Employing the Cauchy-Schwarz inequality, we get the upper bound.
Step 5: Upper bound the regret $\mathcal{R}^{T}$
With all the results in the previous steps, we can eventually prove that the regret $\mathcal{R}^{T}$ can be upper-bounded by $\hat{R}_{h}^{T}$.
5.2 Computational Complexity and Space Complexity
We then prove the computational complexity of employing UCBMQ-GA to train EMDP-GA as below.
Theorem 3.
The computational complexity of employing UCBMQ-GA to train EMDP-GA is of order $\mathcal{O}(H(S+A)T)$, where $H$ is the horizon, $S$ is the size of the state space, $A$ is the size of the action space and $T$ is the number of learning episodes.
The proof is given in Appendix G. Here, we present an intuitive sketch. The updates of $\overline{Q}_{h}^{t}$, $\overline{V}_{h}^{t}$ and $V_{h,s,a}^{t}$ are carried out in an online manner. At time step $h$ of the episode $t$, if $(s_{h}^{t},a_{h}^{t})\neq(s,a)$, both learning rate $\alpha_{H}^{t}(s,a)$ and momentum term $\gamma_{h}^{t}(s,a)$ become zero. Thus, we do not need to update the functions on such $(s,a)$ pairs. Furthermore, the expansions of $Q^{t-1}_{h}$, $\overline{V}^{t-1}_{h}$, and $V^{t-1}_{h,s,a}$ in each episode depend on their corresponding averaged values on $\mathcal{S}_{t-1}$.
In parallel, we have the following result for space complexity.
Theorem 4.
The space complexity of employing UCBMQ-GA to train EMDP-GA is $\mathcal{O}(HS^{2}A)$.
A detailed proof is presented in Appendix G. Intuitively, this lemma holds because we need to store all the bias-value function $V_{h,s,a}^{t}(s^{\prime})$.
Comparison with existing results
As shown in Table 1, the computational complexity of UCBMQ-GA is of the same order as our baseline UCBMQ, smaller than MVP ($\mathcal{O}(HSAT)$) and Bayes-UCBVI ($\mathcal{O}(BHS^{2}AT)$), and a bit larger than UCB-Adv ($\mathcal{O}(HAT)$). The space complexity of UCBMQ-GA is of the same order as UCBMQ, MVP, and Bayes-UCBVI, and still a bit larger than that of UCB-Adv. However, the complexity advances of UCB-Adv are in cost of regret, especially when state space size $S$ and action space size $A$ are large compared with training time $T$, which is the case in an under-explored environment.
6 Conclusion
We mathematically ground the concept of unknown unknowns in reinforcement learning through our proposed episodic Markov decision process with growing awareness (EMDP-GA). By expanding value functions $Q$ and $V$ to newly encountered states with the noninformative value expansion (NIVE) method, our approach effectively addresses the challenge of unknown unknowns. We adapt upper confidence bound momentum Q-learning (UCBMQ) to train the EMDP-GA model at an affordable cost, achieving a regret bound competitive with existing methods not exposed to unknown unknowns.
Applicability
This paper focuses on problems with finite state space and action space. Future works include extending it to continuous state and action spaces.
Appendix A Additional Notations and Preliminaries
Table 2: List of Notation
Description Notation
Size of the state space $S=|\mathcal{S}|$
Size of the action space $A=|\mathcal{A}|$
Number of learning episodes $T$
Horizon (length of an episode) $H$
Aware domain $S_{t},\ t\in[T]$
immediate reward $r_{h}$
Exploration threshold $\zeta=log(\frac{96eHSA(2T+1)}{\delta})$
Indicator function for the occurrence of $(s,a)$ $\chi_{h}^{t}(s,a)$
Visitation count for $(s,a)$ $n_{h}^{t}(s,a)$
First aware domain including $s$ $t(s)$
Policy at episode $t$ $\pi^{t}$
Optimal policy $\pi^{*}$
Optimal Q-value function $Q_{h}^{*}(s,a),\ s\in\mathcal{S},a\in\mathcal{A}$
Optimal value function $V_{h}^{*}(s),\ s\in\mathcal{S}$
Dirac distribution concentrated at $(s_{h+1}^{t})$ $p_{h}^{t}(s^{\prime}|s,a),\ s,s^{\prime}\in\mathcal{S}_{t}$ $\mathring{p}_{h}^{t}(s^{\prime}|s,a),\ s,s^{\prime}\in\mathcal{S}$
Estimated Q-value function $Q_{h}^{t}(s,a),\ s\in\mathcal{S}_{t},\ a\in\mathcal{A}$ $\tilde{Q}_{h}^{t}(s,a),\ s\in\mathcal{S}_{t+1},\ a\in\mathcal{A}$ $\mathring{Q}_{h}^{t}(s,a),\ s\in\mathcal{S},\ a\in\mathcal{A}$
Bias-value function $V_{h,s,a}^{t}(s^{\prime}),\ s,s^{\prime}\in\mathcal{S}_{t}$ $\tilde{V}_{h,s,a}^{t}(s^{\prime}),\ s,s^{\prime}\in\mathcal{S}_{t+1}$ $\mathring{V}_{h,s,a}^{t}(s^{\prime}),\ s,s^{\prime}\in\mathcal{S}$
Upper bound on Q-value function $\overline{Q}_{h}^{t}(s,a),\ s\in\mathcal{S}_{t},\ a\in\mathcal{A}$ $\tilde{\overline{Q}}_{h}^{t}(s,a),\ s\in\mathcal{S}_{t+1},\ a\in\mathcal{A}$ $\mathring{\overline{Q}}_{h}^{t}(s,a),\ s\in\mathcal{S},\ a\in\mathcal{A}$
Upper bound on value function $\overline{V}_{h}^{t}(s),\ s\in\mathcal{S}_{t}$ $\tilde{\overline{V}}_{h}^{t}(s),\ s\in\mathcal{S}_{t+1}$ $\mathring{\overline{V}}_{h}^{t}(s),\ s\in\mathcal{S}$
For $t\in[T]$, we expand $p_{h}^{t}(s^{\prime}|s,a)$, $Q_{h}^{t}(s,a)$, $\overline{V}_{h}^{t}(s)$, $V_{h,s,a}^{t}(s)$, $\overline{Q}_{h}^{t}(s,a)$ to the whole state space $\mathcal{S}$.
$\mathring{p}_{h}^{t}(s^{\prime}|s,a)=\begin{cases}p_{h}^{t}(s^{\prime}|s,a),&s,s^{\prime}\in\mathcal{S}_{t},\\ 0,&s\in\mathcal{S}_{t},s^{\prime}\in\mathcal{S}-\mathcal{S}_{t},\\ p_{h}(s^{\prime}|s,a),&s\in\mathcal{S}-\mathcal{S}_{t},\end{cases}$
$\mathring{Q}_{h}^{t}(s,a)=\begin{cases}Q_{h}^{t}(s,a),&s\in\mathcal{S}_{t},\\ Q_{h,avg}^{t}(a),&s\in\mathcal{S}-\mathcal{S}_{t},\end{cases}$
$\mathring{\overline{V}}_{h}^{t}(s)=\begin{cases}\overline{V}^{t}_{h}(s),&s\in\mathcal{S}_{t},\\ \overline{V}_{h,avg}^{t},&s\in\mathcal{S}-\mathcal{S}_{t},\end{cases}$
$\mathring{V}^{t}_{h,s,a}(s^{\prime})=\begin{cases}V^{t}_{h,s,a}(s^{\prime}),&s,s^{\prime}\in\mathcal{S}_{t},\\ V_{h,s,a,avg}^{t},&s\in\mathcal{S}_{t},s^{\prime}\in\mathcal{S}-\mathcal{S}_{t},\\ V_{h,a,avg}^{t}(s^{\prime}),&s\in\mathcal{S}-\mathcal{S}_{t},s^{\prime}\in\mathcal{S}_{t},\\ V_{h,a,avg}^{t},&s,s^{\prime}\in\mathcal{S}-\mathcal{S}_{t},\end{cases}$
$\mathring{\overline{Q}}_{h}^{t}(s,a)=\begin{cases}Q_{h}^{t}(s,a)+\beta_{h}^{t}(s,a),&n_{h}^{t}(s,a)>0\ \text{(which implies $s\in\mathcal{S}_{t}$)},\\ \mathring{Q}_{h}^{t}(s,a)+H,&n_{h}^{t}(s,a)=0\ \text{or}\ s\in\mathcal{S}-\mathcal{S}_{t}.\end{cases}$
From the definitions, it is easy to see that for any $s,s^{\prime}$ in the aware domain $\mathcal{S}_{t+1}$, we have
$\mathring{Q}_{h}^{t}(s,a)=\tilde{Q}_{h}^{t}(s,a),\ \mathring{\overline{V}}_{h}^{t}(s)=\tilde{\overline{V}}_{h}^{t}(s),\ \mathring{V}^{t}_{h,s,a}(s^{\prime})=\tilde{V}^{t}_{h,s,a}(s^{\prime}),$
and for any state $s\in\mathcal{S}_{t}$,
$\mathring{\overline{Q}}_{h}^{t}(s,a)=\overline{Q}_{h}^{t}(s,a).$
When the count of a state-action pair $n_{h}^{t}(s,a)>0$, we can obtain explicit formulas for the estimate of the Q-function:
$n_{h}^{t}(s,a)Q_{h}^{t}(s,a)=$ $\chi_{h}^{t}(s,a)[r_{h}(s,a)+p_{h}^{t}\tilde{\overline{V}}_{h+1}^{t-1}(s,a)+\mathring{\gamma}_{h}^{t}(s,a)p_{h}^{t}(\tilde{\overline{V}}_{h+1}^{t-1}-\tilde{V}_{h,s,a}^{t-1})(s,a)]$
$+(n_{h}^{t}(s,a)-\chi_{h}^{t}(s,a))\tilde{Q}_{h}^{t-1}(s,a),$
where $\mathring{\gamma}_{h}^{t}(s,a)=H\frac{n_{h}^{t}(s,a)-1}{n_{h}^{t}(s,a)+H}$.
For $s\in\mathcal{S}_{t}-\mathcal{S}_{t-1}$, which means $t(s)=t$, we have $n_{h}^{t}(s,a)=\chi_{h}^{t}(s,a)=1$, hence $\mathring{\gamma}_{h}^{t}(s,a)=0$. Therefore, the following holds
$Q_{h}^{t}(s,a)=r_{h}(s,a)+p_{h}^{t}\tilde{\overline{V}}_{h+1}^{t-1}(s,a).$
For $s\in\mathcal{S}_{t-1}$, we have
$n_{h}^{t}(s,a)Q_{h}^{t}(s,a)=$ $\chi_{h}^{t}(s,a)[r_{h}(s,a)+p_{h}^{t}\tilde{\overline{V}}_{h+1}^{t-1}(s,a)+\mathring{\gamma}_{h}^{t}(s,a)p_{h}^{t}(\tilde{\overline{V}}_{h+1}^{t-1}-\tilde{V}_{h,s,a}^{t-1})(s,a)]$
$+n_{h}^{t-1}(s,a)Q_{h}^{t-1}(s,a),$
$Q_{h}^{t}(s,a)$ $=r_{h}(s,a)+\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[p_{h}^{k}\tilde{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)p_{h}^{k}(\tilde{\overline{V}}_{h+1}^{k-1}-\tilde{V}_{h,s,a}^{k-1})(s,a)].$
From all above, we can obtain explicit formulas for $Q_{h}^{t}$,
$Q_{h}^{t}(s,a)=r_{h}(s,a)+\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)].$
We can do the same with the bias value function when $n_{h}^{t}(s,a)>0$. For $s,s^{\prime}\in\mathcal{S}_{t}$,
$\mathring{V}_{h,s,a}^{t}(s^{\prime})$ $=V_{h,s,a}^{t}(s^{\prime})$
$=\eta_{h}^{t}(s,a)\tilde{\overline{V}}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))\tilde{V}_{h,s,a}^{t-1}(s^{\prime})$
$=\eta_{h}^{t}(s,a)\mathring{\overline{V}}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))\mathring{V}_{h,s,a}^{t-1}(s^{\prime}).$
For $s\in\mathcal{S},s^{\prime}\in\mathcal{S}-\mathcal{S}_{t}$,
$\mathring{V}_{h,s,a}^{t}(s^{\prime})$ $=V_{h,s,a,avg}^{t}$
$=\frac{1}{|\mathcal{S}_{t}|}\sum_{s^{\prime}\in\mathcal{S}_{t}}V_{h,s,a}^{t}(s^{\prime})$
$=\frac{1}{|\mathcal{S}_{t}|}\sum_{s^{\prime}\in\mathcal{S}_{t}}\eta_{h}^{t}(s,a)\tilde{\overline{V}}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))\tilde{V}_{h,s,a}^{t-1}(s^{\prime})$
$=\eta_{h}^{t}(s,a)\mathring{\overline{V}}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))\mathring{V}_{h,s,a}^{t-1}(s^{\prime}).$
For $s\in\mathcal{S}-\mathcal{S}_{t}$, we get $n_{h}^{t}(s,a)=0$.
We can conclude
$\mathring{V}_{h,s,a}^{t}(s^{\prime})$ $=\eta_{h}^{t}(s,a)\mathring{\overline{V}}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))\mathring{V}_{h,s,a}^{t-1}(s^{\prime})$
$=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[\mathring{\overline{V}}_{h+1}^{k-1}(s^{\prime})+\mathring{\gamma}_{h}^{k}(s,a)(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s^{\prime})]$ (1)
$=\sum_{k=t(s)}^{t}\tilde{\eta}_{h}^{t,k}(s,a)\mathring{\overline{V}}_{h+1}^{k-1}(s^{\prime}),$ (2)
where $\tilde{\eta}_{h}^{t,k}(s,a)=\eta_{h}^{k}(s,a)\prod\limits_{l=k+1}^{t}(1-\eta_{h}^{l}(s,a))$.
We prove the following lemma to have a basic knowledge about $\overline{V}^{t}_{h}(s)$ and $V^{t}_{h,s,a}(s^{\prime})$.
Lemma 2.
For $s,s^{\prime}\in\mathcal{S}_{t}$, the following holds almost surely:
  • the sequence $(\overline{V}^{t}_{h}(s))_{t\geq t(s)}$ is non-increasing,
  • $0\leq\overline{V}_{h}^{t}(s)\leq H$,
  • $\overline{V}^{t}_{h+1}(s^{\prime})\leq V^{t}_{h,s,a}(s^{\prime})\leq H$.
When $t\geq t(s)$, we have $\tilde{\overline{V}}_{h}^{t}=\overline{V}_{h}^{t}$.
We can get from the construction of $\overline{V}_{h}^{t}(s)$:
$\overline{V}_{h}^{t}(s)=clip(\mathop{max}\limits_{a\in\mathcal{A}}\overline{Q}_{h}^{t}(s,a),0,\tilde{\overline{V}}_{h}^{t-1}(s))$
that
$\tilde{\overline{V}}_{h}^{t(s)-1}(s)\geq\overline{V}_{h}^{t(s)}(s)\geq\overline{V}_{h}^{t(s)+1}(s)\geq\cdots\geq 0.$
To prove that $\overline{V}_{h}^{t}(s)\leq H,s\in\mathcal{S}_{t}$, we proceed by induction. We have $\overline{V}_{h}^{0}=H$, hence this claim holds when $t=0$. Assuming that $\overline{V}_{h}^{t-1}(s)\leq H,s\in\mathcal{S}_{t-1}$. For $s\in\mathcal{S}_{t-1}\subset\mathcal{S}_{t}$, we have
$\overline{V}_{h}^{t}(s)\leq\overline{V}_{h}^{t-1}(s)\leq H.$
For $s\in\mathcal{S}_{t}-\mathcal{S}_{t-1}$, we have
$\overline{V}_{h}^{t}(s)\leq\tilde{\overline{V}}_{h}^{t-1}(s)=\frac{1}{|\mathcal{S}_{t-1}|}\sum_{\hat{s}\in\mathcal{S}_{t-1}}\overline{V}_{h}^{t-1}(\hat{s})\leq H.$
To prove the third claim, we proceed by induction. When $t=0$, the claims holds. Assuming $\overline{V}^{t-1}_{h+1}(s^{\prime})\leq V^{t-1}_{h,s,a}(s^{\prime})\leq H$ for $s,s^{\prime}\in\mathcal{S}_{t-1}$. We consider two different cases of $s\in\mathcal{S}_{t}$.
Case 1: $s\in\mathcal{S}_{t}-\mathcal{S}_{t-1}$, which means $t(s)=t$ and $\chi_{h}^{t}(s,a)=n_{h}^{t}(s,a)\leq 1$. Since
$\eta_{h}^{t}(s,a)=\chi_{h}^{t}(s,a)\frac{H}{H+n_{h}^{t}(s,a)}\frac{n_{h}^{t}(s,a)-1}{n_{h}^{t}(s,a)}+\chi_{h}^{t}(s,a)\frac{1}{n_{h}^{t}(s,a)}=0,$
the algorithm updates the bias-value function $V_{h,s,a}^{t}(s^{\prime})=\tilde{V}_{h,s,a}^{t-1}(s^{\prime})$.
For $s^{\prime}\in\mathcal{S}_{t-1}$, we have
$V_{h,s,a}^{t}(s^{\prime})=\tilde{V}_{h,s,a}^{t-1}(s^{\prime})=\frac{1}{|\mathcal{S}_{t-1}|}\sum_{\hat{s}\in\mathcal{S}_{t-1}}{V_{h,\hat{s},a}^{t-1}}(s^{\prime})\geq\overline{V}_{h+1}^{t-1}(s^{\prime})\geq\overline{V}_{h+1}^{t}(s^{\prime}).$
For $s^{\prime}\in\mathcal{S}_{t}-\mathcal{S}_{t-1}$, which means $t=t(s^{\prime})$, we have
$V_{h,s,a}^{t}(s^{\prime})$ $=\tilde{V}_{h,s,a}^{t-1}(s^{\prime})$
$=\frac{1}{|\mathcal{S}_{t-1}|^{2}}\sum_{\hat{s},\hat{s^{\prime}}\in\mathcal{S}_{t-1}}{V_{h,\hat{s},a}^{t-1}}(\hat{s^{\prime}})$
$\geq\frac{1}{|\mathcal{S}_{t-1}|}\sum_{\hat{s}\in\mathcal{S}_{t-1}}{\overline{V}_{h+1}^{t-1}(\hat{s})}$
$=\tilde{\overline{V}}_{h+1}^{t}(s^{\prime})\geq\overline{V}_{h+1}^{t}(s^{\prime}).$
Case 2: $s\in\mathcal{S}_{t-1}$, the algorithm updates the bias-value function
$V_{h,s,a}^{t}(s^{\prime})=\eta_{h}^{t}(s,a)\tilde{\overline{V}}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))\tilde{V}_{h,s,a}^{t-1}(s^{\prime}).$
Then we consider the cases of $s^{\prime}\in\mathcal{S}_{t-1}$ and $s^{\prime}\in\mathcal{S}_{t}-\mathcal{S}_{t-1}$. For $s^{\prime}\in\mathcal{S}_{t-1}$, we have
$V_{h,s,a}^{t}(s^{\prime})$ $=\eta_{h}^{t}(s,a)\tilde{\overline{V}}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))\tilde{V}_{h,s,a}^{t-1}(s^{\prime})$
$=\eta_{h}^{t}(s,a)\overline{V}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))V_{h,s,a}^{t-1}(s^{\prime})$
$\geq\eta_{h}^{t}(s,a)\overline{V}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))\overline{V}_{h+1}^{t-1}(s^{\prime})$
$=\overline{V}_{h+1}^{t-1}(s^{\prime})\geq\overline{V}_{h+1}^{t}(s^{\prime}).$
For $s^{\prime}\in\mathcal{S}_{t}-\mathcal{S}_{t-1}$, we have
$V_{h,s,a}^{t}(s^{\prime})$ $=\eta_{h}^{t}(s,a)\tilde{\overline{V}}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))\tilde{V}_{h,s,a}^{t-1}(s^{\prime})$
$=\eta_{h}^{t}(s,a)\frac{1}{|\mathcal{S}_{t-1}|}\sum_{\hat{s}\in\mathcal{S}_{t-1}}\overline{V}_{h+1}^{t-1}(\hat{s})+(1-\eta_{h}^{t}(s,a))\frac{1}{|\mathcal{S}_{t-1}|}\sum_{\hat{s}\in\mathcal{S}_{t-1}}V_{h,s,a}^{t-1}(\hat{s})$
$\geq\eta_{h}^{t}(s,a)\frac{1}{|\mathcal{S}_{t-1}|}\sum_{\hat{s}\in\mathcal{S}_{t-1}}\overline{V}_{h+1}^{t-1}(\hat{s})+(1-\eta_{h}^{t}(s,a))\frac{1}{|\mathcal{S}_{t-1}|}\sum_{\hat{s}\in\mathcal{S}_{t-1}}\overline{V}_{h+1}^{t-1}(\hat{s})$
$=\tilde{\overline{V}}_{h+1}^{t}(s^{\prime})\geq\overline{V}_{h+1}^{t}(s^{\prime}).$
Appendix B A Brief Introduction to Upper Confidence Bound Momentum Q-learning
This Appendix gives an overview of the procedure of UCBMQ. The notations $\chi_{h}^{t}(s,a)$, $n_{h}^{t}(s,a)$, $\alpha_{h}^{t}(s,a)$ and $\gamma_{h}^{t}(s,a)$ are the same in UCBMQ and UCBMQ-GA. For all $(s,a,h,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times[H]\times\mathcal{S}$, UCBMQ is initialized as follows:
$\overline{V}_{h}^{0}(s)=H,\ V_{h,s,a}^{0}(s^{\prime})=H,\ \overline{V}_{H+1}^{t}(s)=0,\ Q_{h}^{0}(s,a)=0.$
In episode $t$, the agent follows a greedy policy based on $\overline{Q}_{h}^{t}(s,a)$ to obtain an episode of length $H$. Next, $Q_{h}^{t}(s,a)$, $V_{h,s,a}^{t}(s^{\prime})$, $\overline{Q}_{h}^{t}(s,a)$ and $\overline{V}_{h}^{t}(s)$ are updated using the data collected in this episode. The algorithm terminates after $T$ episodes.
The updates in UCBMQ described as follows. $Q_{h}^{t}(s,a)$ is updated using the momentum term $\gamma_{h}^{t}(s,a)$:
$Q_{h}^{t}(s,a)=$ $\alpha_{h}^{t}(s,a)[r_{h}(s,a)+p_{h}^{t}\overline{V}_{h+1}^{t-1}(s,a)]$
$+\gamma_{h}^{t}(s,a)p_{h}^{t}(\overline{V}_{h+1}^{t-1}-V_{h,s,a}^{t-1})(s,a)+(1-\alpha_{h}^{t}(s,a))Q_{h}^{t-1}(s,a),$
where $p_{h}^{t}$ denotes the Dirac distribution concentrated at $(s_{h+1}^{t})$ and, for any transition function $p$ and any function $f:\mathcal{S}\rightarrow\mathbb{R}$, we define
$pf(s,a)=\mathbb{E}_{s^{\prime}\sim p(\cdot|s,a)}[f(s^{\prime})].$
$V_{h,s,a}^{t}(s^{\prime})$ is updated according to
$V_{h,s,a}^{t}(s^{\prime})=\eta_{h}^{t}(s,a)\overline{V}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))V_{h,s,a}^{t-1}(s^{\prime}),$
where
$\eta_{h}^{t}(s,a)=\alpha_{h}^{t}(s,a)+\gamma_{h}^{t}(s,a).$
$\overline{Q}_{h}^{t}(s,a)$ is used to upper-bound the optimal Q-value function, which is defined as
$\overline{Q}_{h}^{t}(s,a)=Q_{h}^{t}(s,a)+\beta_{h}^{t}(s,a),$
where $\beta_{h}^{t}(s,a)$ denotes the bonus term. In particular, if $n_{h}^{t}(s,a)=0$ then $\beta_{h}^{t}(s,a)=H$; otherwise
$\beta_{h}^{t}(s,a)=2\sqrt{\frac{\zeta W_{h}^{t}(s,a)}{n_{h}^{t}(s,a)}}+53H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}+\sum_{k=1}^{t}\frac{\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)}{Hlog(T)n_{h}^{t}(s,a)}p_{h}^{k}(V_{h,s,a}^{k-1}-\overline{V}_{h+1}^{k-1})(s,a).$
Here,
$\mathring{\gamma}_{h}^{t}(s,a)=H\frac{n_{h}^{t}(s,a)-1}{n_{h}^{t}(s,a)+H}.$
$\zeta$ is an exploration threshold and $W_{h}^{t}(s,a)$ is a proxy for the variance term defined as
$W_{h}^{t}(s,a)=\sum_{k=1}^{t}\frac{\chi_{h}^{k}(s,a)}{n_{h}^{k}(s,a)}p_{h}^{k}(\overline{V}_{h+1}^{k-1}-\sum_{l=1}^{t}\frac{\chi_{h}^{l}(s,a)}{n_{h}^{l}(s,a)}p_{h}^{l}\overline{V}_{h+1}^{l-1})^{2}(s,a).$
$\overline{V}_{h}^{t}(s)$ is used to upper-bound the optimal value function, which is updated as
$\overline{V}_{h}^{t}(s)=clip(\mathop{\max}_{a\in\mathcal{A}}\overline{Q}_{h}^{t}(s,a),0,\overline{V}_{h}^{t-1}(s)).$
Appendix C Algorithm Chart of Upper Confidence Bound Momentum Q-learning with Growing Awareness
Algorithm 1 UCBMQ-GA for Training EMDA-GA.
1:  Initialise: For all $(s,a,h)$, where $s\in\mathcal{S}_{0}\subset\mathcal{S}$, $V^{0}_{h,s,a}=\overline{V}^{0}_{h}=H$, $Q^{0}_{h}=0$
2:  for $t\in[T]$ do
3:     for $h\in[H]$ do
4:        if $s_{h}^{t}\in\mathcal{S}_{t-1}$ then
5:           Play $a_{h}^{t}\in\mathop{\arg\max}{\overline{Q}_{h}^{t-1}(s_{h}^{t},a)}$
6:        else
7:           Play $a_{h}^{t}\in\mathop{\arg\max}{{Q}_{h,avg}^{t-1}(a)}$
8:        end if
9:        Observe $s_{h+1}^{t}$
10:     end for
11:     $\mathcal{S}_{t}=\mathcal{S}_{t-1}\cup\{s_{1}^{t},\cdots,s_{H}^{t}\}$
12:     Expand $Q^{t-1}_{h}$, $\overline{V}^{t-1}_{h}$ and $V^{t-1}_{h,s,a}$:
$\tilde{Q}^{t-1}_{h}(s,a)=\begin{cases}Q^{t-1}_{h}(s,a),&s\in\mathcal{S}_{t-1},\\ Q^{t-1}_{h,avg}(a),&s\in\mathcal{S}_{t}-\mathcal{S}_{t-1},\end{cases}$
$\tilde{\overline{V}}^{t-1}_{h}(s)=\begin{cases}\overline{V}^{t-1}_{h}(s),&s\in\mathcal{S}_{t-1},\\ \overline{V}_{h,avg}^{t-1},&s\in\mathcal{S}_{t}-\mathcal{S}_{t-1},\end{cases}$
$\tilde{V}^{t-1}_{h,s,a}(s^{\prime})=\begin{cases}V^{t-1}_{h,s,a}(s^{\prime}),&s,s^{\prime}\in\mathcal{S}_{t-1},\\ V_{h,s,a,avg}^{t-1},&s\in\mathcal{S}_{t-1},s^{\prime}\in\mathcal{S}_{t}-\mathcal{S}_{t-1},\\ V_{h,a,avg}^{t-1}(s^{\prime}),&s\in\mathcal{S}_{t}-\mathcal{S}_{t-1},s^{\prime}\in\mathcal{S}_{t-1},\\ V_{h,a,avg}^{t-1},&s,s^{\prime}\in\mathcal{S}_{t}-\mathcal{S}_{t-1}.\end{cases}$
13:     for all $s\in S_{t},a,h$ do
14:        $Q_{h}^{t}(s,a)=\alpha_{h}^{t}(s,a)(r_{h}(s,a)+p_{h}^{t}\tilde{\overline{V}}_{h+1}^{t-1}(s,a))+\gamma_{h}^{t}(s,a)p_{h}^{t}(\tilde{\overline{V}}_{h+1}^{t-1}-\tilde{V}_{h,s,a}^{t-1})(s,a)+(1-\alpha_{h}^{t}(s,a))\tilde{Q}_{h}^{t-1}(s,a)$
15:        $V_{h,s,a}^{t}(s^{\prime})=\eta_{h}^{t}(s,a)\tilde{\overline{V}}_{h+1}^{t-1}(s^{\prime})+(1-\eta_{h}^{t}(s,a))\tilde{V}_{h,s,a}^{t-1}(s^{\prime})$
16:        $\overline{Q}_{h}^{t}(s,a)=Q_{h}^{t}(s,a)+\beta_{h}^{t}(s,a)$
17:        $\overline{V}_{h}^{t}(s)=clip(\mathop{max}\limits_{a\in\mathcal{A}}\overline{Q}_{h}^{t}(s,a),0,\tilde{\overline{V}}_{h}^{t-1}(s))$
18:     end for
19:  end for
Appendix D When a New Observation Hurts Awareness Confidence – A Reverse Bayesian Perspective
Encountering an unknown unknown clearly tells an agent her currently understanding to the environment is problematic. If the agent were a human, this surprise should hurt severely her confidence. We want to model this scenario in agent, and below we show our approach exactly has this property.
We first define a awareness confidence below, as a quantitative measure of a reinforcement learning agent’s confidence to her understanding to the environment, which underpins her confidence to her decisions.
Definition 4 (awareness confidence of an agent).
Suppose an agent has an estimator to the hidden optimal value function $V^{*}$ over a subset of the state space $\mathcal{S}$, defined as function $F:\mathcal{S}^{\prime}\rightarrow\mathbb{R}$, where $\mathcal{S}^{\prime}\subset\mathcal{S}$. The awareness confidence of this agent is defined as the average absolute difference between $F$ and $V^{*}$ over $\mathcal{S}^{\prime}$:
$AC(F)=-\frac{1}{|\mathcal{S}^{\prime}|}\sum_{s\in\mathcal{S}^{\prime}}|F(s)-V^{*}(s)|.$
Remark 6.
The awareness confidence quantifies the potential for improving the estimator. By definition, low awareness confidence indicates that the estimator deviates substantially from the optimal value function, implying that the agent should remain not confident of its estimators. Conversely, high awareness confidence indicates that the estimator closely approximates the optimal value function, thereby justifying the agent’s confidence.
We now prove that the awareness confidence drastically increases when an unknown unknown is encountered.
Theorem 5.
For $t\in[T]$ and $h\in[H]$, where $T$ is the number of learning episodes and $H$ is the horizon, with probability at least $1-\frac{\delta}{2}$, the awareness confidence associated with $\overline{V}_{h}^{t}$ is lower than that associated with $\tilde{\overline{V}}_{h}^{t}$:
$AC(\overline{V}_{h}^{t})<AC(\tilde{\overline{V}}_{h}^{t}).$
Proof.
From Lemma 9, we have
$AC(\tilde{\overline{V}}_{h}^{t})=\frac{1}{|\mathcal{S}_{t+1}|}\sum_{s\in\mathcal{S}_{t+1}}|\tilde{\overline{V}}_{h}^{t}(s)-V^{*}(s)|=\frac{1}{|\mathcal{S}_{t+1}|}\sum_{s\in\mathcal{S}_{t+1}}\tilde{\overline{V}}_{h}^{t}(s)-V^{*}(s),$
$AC(\overline{V}_{h}^{t})=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}|\overline{V}_{h}^{t}(s)-V^{*}(s)|=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}\overline{V}_{h}^{t}(s)-V^{*}(s).$
With Homeland condition in Section 4.2, we have
$\forall s\in\mathcal{S}_{t+1}-\mathcal{S}_{t},V^{*}_{h}(s)\leq\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}V^{*}_{h}(\hat{s})\leq\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}\overline{V}_{h}^{t}(\hat{s})=\tilde{\overline{V}}_{h}^{t}(s).$
Then we get
$\frac{1}{|\mathcal{S}_{t+1}-\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t+1}-\mathcal{S}_{t}}\tilde{\overline{V}}_{h}^{t}(\hat{s})-V^{*}_{h}(\hat{s})$ $\geq\frac{1}{|\mathcal{S}_{t+1}-\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t+1}-\mathcal{S}_{t}}[\frac{1}{|\mathcal{S}_{t}|}\sum_{s^{\prime}\in\mathcal{S}_{t}}\overline{V}_{h}^{t}(s^{\prime})-V^{*}_{h}(s^{\prime})]$
$=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}\overline{V}_{h}^{t}(s)-V^{*}(s).$
Thus, we can conclude $AC(\tilde{\overline{V}}_{h}^{t})\geq AC(\overline{V}_{h}^{t})$. ∎
Remark 7.
This theorem provides a mathematical characterisation to our intuition mentioned in the beginning of this section. In the view of the agent, the environment becomes larger but the agent’s knowledge about the environment is not able to increase at the same time. The knowledge acquired is thus diluted, leading to a degradation in average estimation accuracy, reflecting in the awareness confidence.
Appendix E Proof of Corollary 1
Proof of Corollary 1.
Here we give a special case where $d\overline{V}_{h,avg}^{t}\ (d<1)$ no longer upper-bounds the optimal function on a newly discovered state. Consider an EMDP with a fixed immediate reward function: $\forall(s,a,)\in\mathcal{S}\times\mathcal{A},r_{h}(s,a)=1$. It is easy to get that for every state $s\in\mathcal{S}$, the optimal value function is $V_{h}^{*}(s)=H$. According to Lemma 2, we have $\overline{V}_{h,avg}^{t}\leq H$ for all $h\in[H]$. Therefore, for any $d<1$, we get: $\forall s\in\mathcal{S},\ d\overline{V}_{h,avg}^{t}<H=V_{h}^{*}(s)$. ∎
Appendix F Proof of Theorem 1
Proof of Theorem 1.
We prove the lemma under Assumption 1. For $s^{\prime}\in\mathcal{S}_{0}$, we have $t(s^{\prime})=0\leq\sqrt{T}$. For $s^{\prime}\in\mathcal{S}-\mathcal{S}_{0}$, we have $\forall(s,a)\in\mathcal{S}\times\mathcal{A},P(s^{\prime}|s,a)\geq 1-(\frac{\delta}{3})^{\frac{1}{H[\sqrt{T}]}}\triangleq\theta$. Thus, the following holds $\mathbb{P}(s^{\prime}\notin O_{[\sqrt{T}]})\leq(1-\theta)^{H[\sqrt{T}]}$. From the relationship between $O_{t}$ and $\mathcal{S}_{t}$, we get
$\mathbb{P}(s^{\prime}\notin\mathcal{S}_{[\sqrt{T}]})\leq(1-\theta)^{H[\sqrt{T}]}=\frac{\delta}{3}.$
Finally, we get $t(s^{\prime})\leq[\sqrt{T}]\leq\sqrt{T}$ with probability at least $1-\frac{\delta}{3}$. ∎
Appendix G Proof of Lemma 3 and Lemma 4
Proof of Lemma 3.
We prove the computational complexity of UCBMQ-GA by analysing the cost of each time-step. The computational complexity of UCBMQ-GA per time step comprises function updates and average-value maintenance, which can be performed concurrently. At time-step $h$ of an episode $t$, we only need to update the functions concerning $(s_{h}^{t},a_{h}^{t})$ since for other $(s,a)$, $\alpha_{h}^{t}(s,a)=0$ and $\gamma_{h}^{t}(s,a)=0$. Thus, updating ${Q}_{h}^{t}$ and $\overline{Q}_{h}^{t}$ takes $\mathcal{O}(1)$ time Updating the bias-value function $V_{h,s,a}^{t}$ takes $\mathcal{O}(S)$ time because the update is performing linear interpolation over a vector of length $S$. Updating $\overline{V}_{h}^{t}$ on $s_{h}^{t}$ involves taking the maximum over all $A$, which takes $\mathcal{O}(A)$ time. We can calculate average values every time the functions are updated, and the calculation takes $\mathcal{O}(1)$ time. Hence, each step of updating functions costs $\mathcal{O}(S+A)$, each step of calculating average values costs $\mathcal{O}(1)$ and the total cost is $\mathcal{O}HT(S+A)$. ∎
Proof of Lemma 4.
UCBMQ-GA needs space to store the functions. ${Q}_{h}^{t}$ and $\overline{Q}_{h}^{t}$ both require $\mathcal{O}(SAH)$ space. $\overline{V}_{h}^{t}$ needs $\mathcal{O}(HS)$ space, while $V_{h,s,a}^{t}$ needs $\mathcal{O}(HS^{2}A)$ space. So the space complexity is $\mathcal{O}(HS^{2}A)$. ∎
Appendix H Concentration Events
We firstly introduce the Bernstein-type concentration inequality.
Theorem 6 (Bernstein-type concentration inequality [9]).
Let $(Y_{t})_{t\in\mathbb{N}^{*}},(\omega_{t})_{t\in\mathbb{N}^{*}}$ be two sequences of random variables adapted to a filtration $(\mathcal{F}_{t})_{t\in\mathbb{N}}$. We assume that the weights are in the unit interval $\omega_{t}\in[0,1]$ and predictable, i.e. $\mathcal{F}_{t-1}$ measurable. We also assume that the random variables $Y_{t}$ are bounded $|Y_{t}|\leq b$ and centered $\mathbb{E}[Y_{t}|\mathcal{F}_{t-1}]=0$. Consider the following quantities
$S_{t}\triangleq\sum_{s=1}^{t}\omega_{s}Y_{s},\ V_{t}\triangleq\sum_{s=1}^{t}\omega_{s}^{2}\mathbb{E}[Y_{t}^{2}|\mathcal{F}_{s-1}],\ W_{t}\triangleq\sum_{s=1}^{t}\omega_{s},$
and let $h(s)\triangleq(x+1)log(x+1)-x$ be the Crámer transform of a Poisson distribution of parameter $1$. For all $\delta>0$,
$\mathbb{P}[\exists t\geq 1,(\frac{V_{t}}{b^{2}})h(\frac{b|S_{t}|}{V_{t}+b^{2}})\geq log(\frac{1}{\delta})+log(4e(2t+1))]<\delta.$
The previous inequality can be weakened to obtain a more explicit bound: if $b\geq 1$ with probability at least $1-\delta$, for all $t\geq 1$,
$|S_{t}|\leq\sqrt{2V_{t}log(\frac{4e(2t+1)}{\delta})}+3blog(\frac{4e(2t+1)}{\delta}).$
We define the favorable events as follows:
[…]
$|\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h}^{k-1})(s,a)}+6H\zeta\},$
[…]
$\leq\sqrt{8H^{2}\zeta\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)}+12H^{2}\zeta\},$
[…]
$|\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)|$
$\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)^{2}Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)}+6H^{2}\zeta\},$
[…]
$\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)}+6H\zeta\}.$
We define $\mathcal{E}=\mathcal{E}^{v_{1}}\cap\mathcal{E}^{v_{2}}\cap\mathcal{E}^{m_{\omega}}\cap\mathcal{E}^{m}$ the intersection of all the events above. $\mathcal{E}$ holds with high probability.
Lemma 3.
For the choice
$\zeta=log(\frac{96e(2T+1)}{\delta}),$
the following holds $\mathbb{P}(\mathcal{E})\geq 1-\frac{\delta}{6}$.
Proof.
From Lemma 2 we can get
[…]
Therefore, for $s\in\mathcal{S}_{k}$
$|(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|\leq 2H,$
$|(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)|\leq 4H^{2},$
$|(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)|\leq 2H.$
For $\mathcal{E}^{v_{2}}$, based on Lemma 16, we have $Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)\leq 2H^{2}Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$. Based on $\zeta=log(\frac{32e(2T+1)}{\delta})$ and Theorem 6, we get
$\mathbb{P}((\mathcal{E}^{v_{1}})^{c})\leq\frac{\delta}{24},\ \ \ \mathbb{P}((\mathcal{E}^{v_{2}})^{c})\leq\frac{\delta}{24},\ \ \ \mathbb{P}((\mathcal{E}^{m_{\omega}})^{c})\leq\frac{\delta}{24},\ \ \ \mathbb{P}((\mathcal{E}^{m})^{c})\leq\frac{\delta}{24}.$
Combine all above, we can conclude $\mathbb{P}(\mathcal{E})\geq 1-\frac{\delta}{6}$. ∎
Lemma 4.
On the event $\mathcal{E}$, $\forall t\in[T],\ \forall h\in[H],\ \forall(s,a)\in\mathcal{S}_{t}\times\mathcal{A}$, the following holds
$\ \ \ \ \ |\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)|$
$\leq\frac{1}{4Hlog(T)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+14H^{3}log(T)\zeta,$
$|\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)|\leq\frac{1}{4}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+14H\zeta,$
and
$|\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)|\leq\frac{1}{4}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+44H^{2}\zeta.$
Proof.
For the first claim, we use the fact that $\mathcal{E}^{m_{\omega}}$ holds and for all $s\in\mathcal{S}_{t},s^{\prime}\in\mathcal{S}$ we have
[…]
and get
$\ \ \ \ \ |\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)|$
$\leq\sqrt{2\zeta H\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)}+6H^{2}\zeta$
$\leq\sqrt{2\zeta H^{2}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)}+6H^{2}\zeta$
$\leq\frac{1}{4Hlog(T)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+14H^{3}log(T)\zeta.$
For the second claim, on the event $\mathcal{E}^{m}$, we have
$\ \ \ \ \ |\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)|$
$\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)}+6H\zeta$
$\leq\sqrt{2\zeta H\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)}+6H\zeta$
$\leq\frac{1}{4}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+14H\zeta.$
For the third claim, we use event $\mathcal{E}^{v_{2}}$:
$\ \ \ \ \ |\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)|$
$\leq\sqrt{8H^{2}\zeta\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)}+12H^{2}\zeta$
$\leq\frac{1}{4}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+44H^{2}\zeta.$
This concludes the proof. ∎
For $s\in\mathcal{S}$, Let $\overline{p}_{h}^{t}(s,a)$ and $\overline{p}_{h}^{t}(s)$ denote the probabilities to reach $(s,a)$ and $s$, respectively, at the time-step $h$ under the policy $\pi^{t}$ in the algorithm. We define the following events:
$\mathcal{G}^{Var}$$\triangleq\{\forall t\in[T],\forall h\in[H],\forall(s,a)\in\mathcal{S}_{t}\times\mathcal{A}:$
$|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)|\leq\sum_{k=t(s)}^{t}\overline{p}_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)+8H^{2}\zeta\},$
$\mathcal{G}^{v_{1}}\triangleq\{\forall t\in[T],\forall h\in[H],\forall(s,a)\in\mathcal{S}_{t}$$\times\mathcal{A}:|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}})(s,a)|$
$\leq\frac{1}{4H}\sum_{k=t(s)}^{t}\overline{p}_{n}^{k}(s,a)p_{h}|\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}}|(s,a)+14H^{2}\zeta\},$
$\mathcal{G}^{v_{2}}$$\triangleq\{\forall t\in[T],\forall h\in[H],\forall s\in\mathcal{S}_{t}:$
$|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s)(\mathring{\overline{V}}_{h}^{k-1}-V_{h}^{\pi^{t}})(s)|\leq\frac{1}{4H}\sum_{k=t(s)}^{t}\overline{p}_{n}^{k}(s)|\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}}|(s)+14H^{2}\zeta\},$
We define $\mathcal{G}=\mathcal{G}^{Var}\cap\mathcal{G}^{v_{1}}\cap\mathcal{G}^{v_{2}}$ the intersection of the events. This event holds with high probability.
Lemma 5.
For the choice
$\zeta=log(\frac{96eHSA(2T+1)}{\delta}),$
the following holds $\mathbb{P}(\mathcal{G})\geq 1-\frac{\delta}{6}$.
Proof.
Based on Theorem 6, with probability at least $1-\frac{\delta}{24}$, $\forall t\in[T],\forall h\in[H],\forall(s,a)\in\mathcal{S}_{t}\times\mathcal{A}$, using the facts that for $X\sim Ber(q)$ the following holds $Var(X)=q(1-q)\leq q$ and $\sqrt{xy}\leq x+y$,
$|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)|$ $\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\overline{p}_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)^{2}}+6H^{2}\zeta$
$\leq\sum_{k=t(s)}^{t}\overline{p}_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)+8H^{2}\zeta.$
So, we have
$\ \ \ \ \ |\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}})(s,a)|$
$\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\overline{p}_{n}^{k}(s,a)p_{h}|\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}}|(s,a)^{2}}+6\zeta H^{2}$
$\leq\frac{1}{4H}\sum_{k=t(s)}^{t}\overline{p}_{n}^{k}(s,a)p_{h}|\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}}|(s,a)+14H^{2}\zeta\}.$
Further,
$|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s)(\mathring{\overline{V}}_{h}^{k-1}-V_{h}^{\pi^{t}})(s)|$ $\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\overline{p}_{n}^{k}(s)|\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}}|(s)^{2}}+6\zeta H^{2}$
$\leq\frac{1}{4H}\sum_{k=t(s)}^{t}\overline{p}_{n}^{k}(s)|\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}}|(s)+14H^{2}\zeta.$
Therefore, we get
$\mathbb{P}((\mathcal{G}^{Var})^{c})\leq\frac{\delta}{24},\ \mathbb{P}((\mathcal{G}^{V_{1}})^{c})\leq\frac{\delta}{24},\ \mathbb{P}((\mathcal{G}^{V_{2}})^{c})\leq\frac{\delta}{24}.$
We can conclude that $\mathbb{P}(\mathcal{G})\geq 1-\frac{\delta}{6}$. ∎
We define the event $\mathcal{D}=\mathcal{E}\cap\mathcal{G}$. Combine Lemma 3 and Lemma 5, we get:
Lemma 6.
For the choice
$\zeta=log(\frac{96eHSA(2T+1)}{\delta}),$
the following holds $\mathbb{P}(\mathcal{D})\geq 1-\frac{\delta}{3}$.
Appendix I Optimism
We use $r_{h}(s,a)+p_{h}\mathring{V}_{h,s,a}^{t}(s,a)$ to estimate $Q_{h}^{t}(s,a)$. The gap between $r_{h}(s,a)+p_{h}\mathring{V}_{h,s,a}^{t}(s,a)$ and $Q_{h}^{t}(s,a)$ proves to be controlled by $\beta_{h}^{t}(s,a)$. Finally, we can use $\mathring{\overline{Q}}_{h}^{t}(s,a)$ to upper-bound $Q_{h}^{*}(s,a)$ and $\mathring{\overline{V}}_{h}^{t}(s)$ to upper-bound $V_{h}^{*}(s,a)$.
Lemma 7.
On the event $\mathcal{E}$, $\forall t\in[T],\ \forall h\in[H],\ \forall(s,a)\in\mathcal{S}\times\mathcal{A}$, if $n_{n}^{t}(s,a)>0$, the following holds,
$\ \ \ \ \ |Q_{h}^{t}(s,a)-r_{h}(s,a)-p_{h}\mathring{V}_{h,s,a}^{t}(s,a)|$
$\leq\sqrt{\frac{2}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h}^{k-1})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}+20H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
$\ \ \ \ \ +\frac{1}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a),$
where $\mathring{\gamma}_{h}^{t}(s,a)=H\frac{n_{h}^{t}(s,a)-1}{n_{h}^{t}(s,a)+H}$.
Proof.
$n_{h}^{t}(s,a)>0$ implies $s\in\mathcal{S}_{t}$. So, we have
$\ \ \ \ \ Q_{h}^{t}(s,a)-r_{h}(s,a)-p_{h}\mathring{V}_{h,s,a}^{t}(s,a)$
$=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)]-p_{h}\mathring{V}_{h,s,a}^{t}(s,a)$
$=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)]$
$\leq|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|$
$\ \ \ \ \ +|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)|.$
We can upper-bound the first term of the right-hand by the definition of $\mathcal{E}$:
$\ \ \ \ \ |\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|$
$\leq\sqrt{\frac{2}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h}^{k-1})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}+6H\frac{\zeta}{n_{h}^{t}(s,a)}.$
For the second term, we use Lemma 4, and get
$\ \ \ \ \ |\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)|$
$\leq\frac{1}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+14H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}.$
This concludes the proof. ∎
We design a bonus to compensate for the bias in the previous lemma.
Lemma 8.
On the event $\mathcal{E}$, $\forall t\in[T],\ \forall h\in[H],\ \forall(s,a)\in\mathcal{S}\times\mathcal{A}$, if $n_{n}^{t}(s,a)>0$, the following holds
$\beta_{h}^{t}(s,a)\geq$ $\sqrt{\frac{2}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}+20H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
$+\frac{1}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{t}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}_{h,s,a}^{k-1}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a).$
Proof.
The definition of bonus is: for $s\in\mathcal{S}_{t}$
$\beta_{h}^{t}(s,a)=$ $2\sqrt{W_{h}^{t}(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}+53H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
$+\frac{1}{Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{V}_{h,s,a}^{k-1}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a),$
where
$W_{h}^{t}(s,a)=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)-(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}.$
Since we do not know the true transitions $p_{h}$, the design of the bonus only needs the observed data.
Based on Lemma 4, we can control the correction term as follows,
$\ \ \ \ \ |\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)|$ (3)
$\leq\frac{1}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+14H^{3}\frac{log(T)\zeta}{n_{h}^{t}(s,a)}.$
As for the proxy of the variance term $W_{h}^{t}(s,a)$, applying Lemma 4 and the definition of $\mathcal{E}$ and $\sqrt{xy}\leq x+y$, we get
$\ \ \ \ \ \frac{1}{n_{h}^{t}(s,a)}|\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)|$ (4)
$\leq\frac{1}{4n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+44H^{2}\frac{\zeta}{n_{h}^{t}(s,a)},$
and
$\ \ \ \ \ |(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}-(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}|$
$\leq|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}+p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|\cdot|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|$
$\leq\frac{2H}{n_{h}^{t}(s,a)}|\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|$
$\leq H\sqrt{8\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}+12H^{2}\frac{\zeta}{n_{h}^{t}(s,a)}$
$\leq\frac{1}{4n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+44H^{2}\frac{\zeta}{n_{h}^{t}(s,a)}.$ (5)
Then, combining eq. (4), eq. (5) and Jensen’s inequality, we get
$W_{h}^{t}(s,a)$ $=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)-(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
$\geq\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)-(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
$\ \ \ \ \ -\frac{1}{2n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)-88H^{2}\frac{\zeta}{n_{h}^{t}(s,a)}$
$\geq\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)-\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(p_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
$\ \ \ \ \ -\frac{1}{2n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)-88H^{2}\frac{\zeta}{n_{h}^{t}(s,a)}$
$=\frac{1}{2n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)-88H^{2}\frac{\zeta}{n_{h}^{t}(s,a)}.$ (6)
Finally, with eq. (3) and eq. (6), we can conclude
$\beta_{h}^{t}(s,a)$ $\geq 2\sqrt{(W_{h}^{t}(s,a)+88H^{2}\frac{\zeta}{n_{h}^{t}(s,a)})\frac{\zeta}{n_{h}^{t}(s,a)}}+(53-2\sqrt{88}-14)H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
$\ \ \ \ \ +\frac{3}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}_{h,s,a}^{k-1}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$
$\geq 2\sqrt{\frac{2}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}+20H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
$\ \ \ \ \ +\frac{1}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}_{h,s,a}^{k-1}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a),$
which also relies on the fact that $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$.
This concludes the proof. ∎
Now we are ready to prove Lemma 9 on the event $\mathcal{E}$ and with the Homeland condition in Section 4.2.
Lemma 9.
The following holds with probability at least $1-\delta$, $\forall t\in[T],\ \forall(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ (also holds for $H+1$ for the value function), we have
$\mathring{\overline{Q}}_{h}^{t}(s,a)\geq Q^{*}_{h}(s,a),\ \ \ \text{and}\ \ \ \mathring{\overline{V}}_{h}^{t}(s)\geq V_{h}^{*}(s).$
Proof.
We proceed by induction on $t$. The two claims hold trivially for $t=0$ as a result of the initialisation of the algorithm. Assume the two claims hold for $k\leq t-1$. We consider the case of $t$. For $h=H+1$, we have
$\mathring{\overline{Q}}_{h}^{t}(s,a)=Q^{*}_{h}(s,a)=0,$
and
$\mathring{\overline{V}}_{h}^{t}(s)=V_{h}^{*}(s)=0.$
We assume
$\mathring{\overline{Q}}_{h+1}^{t}(s,a)\geq Q^{*}_{h+1}(s,a),$
and
$\mathring{\overline{V}}_{h+1}^{t}(s)\geq V_{h+1}^{*}(s).$
If $n_{h}^{t}(s,a)=0$ and $s\in\mathcal{S}_{t}$, we have
$\mathring{\overline{Q}}_{h}^{t}(s,a)$ $=\frac{1}{|\mathcal{S}_{t(s)-1}|}\sum_{\hat{s}\in\mathcal{S}_{t(s)-1}}Q_{h}^{t(s)-1}(\hat{s},a)+H$
$\geq\frac{1}{|\mathcal{S}_{t(s)-1}|}\sum_{\hat{s}\in\mathcal{S}_{t(s)-1}}\overline{Q}_{h}^{t(s)-1}(\hat{s},a)$
$\geq\frac{1}{|\mathcal{S}_{t(s)-1}|}\sum_{\hat{s}\in\mathcal{S}_{t(s)-1}}{Q}_{h}^{*}(\hat{s},a)$
$\geq Q^{*}_{h}(s,a).$
If $s\in\mathcal{S}-\mathcal{S}_{t}$, we have
$\mathring{\overline{Q}}_{h}^{t}(s,a)$ $=\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}Q_{h}^{t}(\hat{s},a)+H$
$\geq\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}\overline{Q}_{h}^{t}(\hat{s},a)$
$\geq\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}{Q}_{h}^{*}(\hat{s},a)$
$\geq Q^{*}_{h}(s,a).$
If $n_{h}^{t}(s,a)>0$, applying Lemmas 7 and 8, and because the event $\mathcal{E}$ holds, we get
$\mathring{\overline{Q}}_{h}^{t}(s,a)$ $=Q_{h}^{t}(s,a)+\beta_{h}^{t}(s,a)$
$\geq r_{h}(s,a)+p_{h}\mathring{V}_{h,s,a}^{t}(s,a)$
$=r_{h}(s,a)+p_{h}(\sum_{k=t(s)}^{t}\tilde{\eta}_{h}^{t,k}(s,a)\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$
$\geq r_{h}(s,a)+p_{h}V^{*}_{h}(s,a)=Q^{*}_{h}(s,a).$
For $s\in\mathcal{S}_{t}$, note that
$\mathring{\overline{V}}_{h}^{t}(s)=\overline{V}_{h}^{t}(s)\geq clip(\mathop{max}\limits_{a\in\mathcal{A}}\overline{Q}_{h}^{t}(s,a),0,H)\geq\mathop{max}\limits_{a\in\mathcal{A}}Q_{h}^{*}(s,a)=V^{*}_{h}(s).$
For $s\in\mathcal{S}-\mathcal{S}_{t}$, we have
$\mathring{\overline{V}}_{h}^{t}(s)$ $=\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}\overline{V}_{h}^{t}(\hat{s})$
$\geq\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}clip(\mathop{max}\limits_{a\in\mathcal{A}}\overline{Q}_{h}^{t}(\hat{s},a),0,H)$
$\geq\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}\mathop{max}\limits_{a\in\mathcal{A}}{Q}_{h}^{*}(\hat{s},a)$
$\geq\mathop{max}\limits_{a\in\mathcal{A}}\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}{Q}_{h}^{*}(\hat{s},a)$
$\geq\mathop{max}\limits_{a\in\mathcal{A}}{Q}_{h}^{*}(s,a)=V^{*}_{h}(s).$
The proof is completed. ∎
Appendix J Proof of Theorem 2
Let $\tilde{n}_{h}^{t}(s,a)=\mathop{max}\{n_{h}^{t}(s,a),1\}$. We first give two important lemmas.
Lemma 10.
On the event $\mathcal{E}$, with Homeland condition in Section 4.2, $\forall t\in[T],\ \forall h\in[H],\ \forall(s,a)\in\mathcal{S}\times\mathcal{A}$, we have,
$\ \ \ \ \ |\mathring{Q}_{h}^{t}(s,a)-r_{h}(s,a)-p_{h}\mathring{V}_{h,s,a}^{t}(s,a)|$
$\leq\sqrt{\frac{4}{\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{\tilde{n}_{h}^{t}(s,a)}}+24H^{3}\frac{\zeta log(T)}{\tilde{n}_{h}^{t}(s,a)}$
$\ \ \ \ \ +\frac{2}{Hlog(T)\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-{V}_{h+1}^{\pi_{k}})(s,a).$
Proof.
we first prove $|\mathring{Q}_{h}^{t}(s,a)|\leq H^{2}$ by induction. When $t=0$, the claim trivially holds. Assume that the claim holds for $k\leq t-1$. For $s\in\mathcal{S}_{t},\ n_{h}^{t}(s,a)=0$, we have
$|Q_{h}^{t}(s,a)|=|\mathring{Q}_{h}^{t}(s,a)|\leq H^{2}.$
For $s\in\mathcal{S}_{t},\ n_{h}^{t}(s,a)>0$, we have
$\ \ \ \ \ |Q_{h}^{t}(s,a)|$
$\leq|r_{h}(s,a)|+|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)]|$
$\leq 1+|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[H+\frac{H^{2}}{2}|\leq H^{2}.$
For $s\in\mathcal{S}-\mathcal{S}_{t}$, we have
$|\mathring{Q}_{h}^{t}(s,a)|=|\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}Q_{h}^{t}(s,a)|\leq H^{2}.$
Therefore, we can conclude $\forall(s,a)\in\mathcal{S}\times\mathcal{A}:\ |\mathring{Q}_{h}^{t}(s,a)|\leq H^{2}$.
For $n_{h}^{t}(s,a)=0$, the bound trivially holds because
$|\mathring{Q}_{h}^{t}(s,a)-r_{h}(s,a)-p_{h}\mathring{V}_{h,s,a}^{t}(s,a)|\leq|\mathring{Q}_{h}^{t}(s,a)|+|r_{h}(s,a)|+|p_{h}\mathring{V}_{h,s,a}^{t}(s,a)|\leq H^{2}+H+1.$
For $n_{h}^{t}(s,a)>0$, we have on the event $\mathcal{E}$ with Lemma 7
$\ \ \ \ \ |\mathring{Q}_{h}^{t}(s,a)-r_{h}(s,a)-p_{h}\mathring{V}_{h,s,a}^{t}(s,a)|$
$\leq\sqrt{\frac{2}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h}^{k-1})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}+20H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
$\ \ \ \ \ +\frac{1}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a).$
Applying $\mathring{V}_{h,s,a}^{t}\geq V^{*}_{h+1}\geq V_{h+1}^{\pi^{k}}$ and eq. (1), we can get
$\ \ \ \ \ \frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}_{h,s,a}^{k-1}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$
$\leq\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-{V}_{h+1}^{\pi^{k}})(s,a).$ (7)
Applying $H\geq\mathring{\overline{V}}_{h}^{k}\geq V^{*}_{h}\geq V_{h}^{\pi^{k+1}}$ and Lemma 16, we have
$\ \ \ \ \ \frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$
$\leq\frac{2}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}({V}_{h+1}^{\pi^{k}})(s,a)+\frac{2H}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a).$
Applying $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ and $\sqrt{xy}\leq x+y$, we get
$\ \ \ \ \ \sqrt{\frac{2}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$
$\leq\sqrt{\frac{4}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}({V}_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$
$\ \ \ \ \ +\sqrt{\frac{4H}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$
$\leq\sqrt{\frac{4}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}({V}_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$
$\ \ \ \ \ +\frac{1}{Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}+4H^{2}\frac{log(T)\zeta}{n_{h}^{t}(s,a)}.$ (8)
Combining eq. (7) and eq. (8), we can conclude
$\ \ \ \ \ |\mathring{Q}_{h}^{t}(s,a)-r_{h}(s,a)-p_{h}\mathring{V}_{h,s,a}^{t}(s,a)|$
$\leq\sqrt{\frac{4}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}+24H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
$\ \ \ \ \ +\frac{2}{Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}^{k-1}_{h+1}-V_{h+1}^{\pi^{k}})(s,a).$
Then, we give an upper bound on the bonus.
Lemma 11.
On the event $\mathcal{E}$, under Homeland condition in Section 4.2, $\forall t\in[T],\ \forall h\in[H],\ \forall(s,a)\in\mathcal{S}\times\mathcal{A}$, the following holds
$\beta_{h}^{t}(s,a)$ $\leq 2\sqrt{\frac{3}{\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{\tilde{n}_{h}^{t}(s,a)}}+106H^{3}\frac{log(T)\zeta}{\tilde{n}_{h}^{t}(s,a)}$
$\ \ \ \ \ +\frac{3}{Hlog(T)\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a).$
Proof.
If $n_{h}^{t}(s,a)=0$, the bound is trivially true because $\beta_{h}^{t}(s,a)=H$. We now consider the case $n_{h}^{t}(s,a)>0$. Using eq. (4), eq. (5) and $H\geq\mathring{\overline{V}}_{h}^{k}\geq V^{*}_{h}$, we obtain
$W_{h}^{t}(s,a)$ $\leq\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)-(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
$\ \ \ \ \ +\frac{1}{2n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+88H^{2}\frac{\zeta}{n_{h}^{t}(s,a)}$
$=Var_{p_{h}}(V_{h+1}^{*})(s,a)+\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}((\mathring{\overline{V}}_{h+1}^{k-1})^{2}-(V_{h+1}^{*})^{2})(s,a)$
$\ \ \ \ \ +(p_{h}V_{H+1}^{*}(s,a))^{2}-(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
$\ \ \ \ \ +\frac{1}{2n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+88H^{2}\frac{\zeta}{n_{h}^{t}(s,a)}$
$\leq Var_{p_{h}}(V_{h+1}^{*})(s,a)+\frac{1}{2n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$
$\ \ \ \ \ \frac{2H}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V^{*}_{h+1})(s,a)+88H^{2}\frac{\zeta}{n_{h}^{t}(s,a)}.$
Applying $H\geq\mathring{\overline{V}}_{h}^{k}\geq V_{h}^{*}\geq V_{h}^{\pi^{k+1}}$ and applying Lemma 16 to the terms $Var_{p_{h}}(V^{*}_{h+1})(s,a)$ and $Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$, we obtain
$W_{h}^{t}(s,a)$ $\leq\frac{3}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)$
$\ \ \ \ \ +\frac{2H}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(V^{*}_{h+1}-V_{h+1}^{\pi^{k}})(s,a)$
$\ \ \ \ \ +\frac{H}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a)$
$\ \ \ \ \ +\frac{2H}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V^{*}_{h+1})(s,a)+88H^{2}\frac{\zeta}{n_{h}^{t}(s,a)}$
$\leq\frac{3}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)$
$\ \ \ \ \ +\frac{5H}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V^{\pi^{k}}_{h+1})(s,a)+88H^{2}\frac{\zeta}{n_{h}^{t}(s,a)}.$ (9)
Combine eq. (9) with $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ and $\sqrt{xy}\leq x+y$ we can upper-bound the variance term of the bonus
$2\sqrt{W_{h}^{t}(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$ $\leq 2\sqrt{\frac{3}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$
$\ \ \ \ \ +\sqrt{\frac{20H}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V^{\pi^{k}}_{h+1})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$
$\ \ \ \ \ +2\sqrt{88}H\frac{\zeta}{n_{h}^{t}(s,a)}$
$\leq 2\sqrt{\frac{3}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$
$\ \ \ \ \ +\frac{1}{Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V^{\pi^{k}}_{h+1})(s,a)+\frac{20H^{2}\zeta}{n_{h}^{t}(s,a)}$
$\ \ \ \ \ +19H\frac{\zeta}{n_{h}^{t}(s,a)}$
$\leq 2\sqrt{\frac{3}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$
$\ \ \ \ \ +\frac{1}{Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V^{\pi^{k}}_{h+1})(s,a)$
$\ \ \ \ \ +\frac{39H^{2}\zeta}{n_{h}^{t}(s,a)}.$ (10)
Applying Lemma 4 and eq. (7), we have
$\ \ \ \ \ \frac{1}{Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{V}_{h,s,a}^{k-1}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$
$\leq 14H^{3}\frac{log(T)\zeta}{n_{h}^{t}(s,a)}+\frac{5}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}_{h,s,a}^{k-1}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$
$\leq 14H^{3}\frac{log(T)\zeta}{n_{h}^{t}(s,a)}+\frac{5}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a).$ (11)
Combining eq. (10) and eq. (11) we obtain the result of this lemma. ∎
Now we are ready to prove Theorem 2.
Proof of Theorem 2.
We assume the event $\mathcal{D}$, Homeland condition in Section 4.2, and Assumption 1 hold.
From Lemma 6, we have that $\mathcal{D}$, Homeland condition in Section 4.2 and Assumption 1 hold with probability at least $1-\delta$. Fix $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $t\in[T]$.
Step 1: Upper-bound $(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)$     Applying Lemmas 10 and 11, we get
$(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)\leq$ $p_{h}(\mathring{V}^{t}_{h,s,a}-V_{h+1}^{\pi^{t+1}})(s,a)+130H^{3}\frac{log(T)\zeta}{\tilde{n}_{h}^{t}(s,a)}$
$+\sqrt{\frac{30}{\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{\tilde{n}_{h}^{t}(s,a)}}$
$+\frac{5}{Hlog(T)\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a).$ (12)
Step 2: Upper-bound the local optimistic regret $\hat{R}_{h}^{T}(s,a)$     We define
$\hat{R}_{h}^{T}(s,a)\triangleq\sum_{t=0}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)=\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a).$
Based on this, eq. (12) yields the follows,
$\hat{R}_{h}^{T}(s,a)\leq$ $\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(\mathring{V}^{t}_{h,s,a}-V_{h+1}^{\pi^{t+1}})(s,a)+130H^{3}\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)\frac{log(T)\zeta}{\tilde{n}_{h}^{t}(s,a)}$
$+\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)\sqrt{\frac{30}{\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{\tilde{n}_{h}^{t}(s,a)}}$
$+\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)\frac{5}{Hlog(T)\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a).$
For the first term, we have the decomposition
$p_{h}(\mathring{V}^{t}_{h,s,a}-V_{h+1}^{\pi^{t+1}})(s,a)=p_{h}(\mathring{V}^{t}_{h,s,a}-V^{*}_{h+1})(s,a)+p_{h}(V^{*}_{h+1}-V_{h+1}^{\pi^{t+1}})(s,a).$ (13)
Then, applying eq. (2) and Lemma 14, we obtain
$\ \ \ \ \ \sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(\mathring{V}^{t}_{h,s,a}-V_{h+1}^{*})(s,a)$
$=\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)\mathbb{I}_{\{n_{h}^{t}(s,a)=0\}}p_{h}(\mathring{V}^{t}_{h,s,a}-V_{h+1}^{*})(s,a)$
$\ \ \ \ \ +\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)\mathbb{I}_{\{n_{h}^{t}(s,a)>0\}}\sum_{k=t(s)}^{t}p_{h}(\mathring{\overline{V}}^{k-1}_{h+1}-V_{h+1}^{*})(s,a)$
$\leq H+\sum_{k=t(s)}^{T-1}(\sum_{t=k}^{T-1}\chi_{h}^{t+1}(s,a)\tilde{\eta}_{h}^{t,k}(s,a))p_{h}(\mathring{\overline{V}}^{k-1}_{h+1}-V_{h+1}^{*})(s,a)$
$\leq H+(1+\frac{1}{H})\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t-1}-V^{*}_{h+1})(s,a).$ (14)
Combining eq. (13), eq. (14), and $V^{*}_{h+1}\geq V_{h+1}^{\pi^{k+1}}$, we get
$\ \ \ \ \ \sum_{t=0}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(\mathring{V}^{t}_{h,s,a}-V_{h+1}^{\pi^{t+1}})(s,a)$
$\leq\sum_{t=0}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(V^{*}_{h+1}-V_{h+1}^{\pi^{t+1}})(s,a)$
$\ \ \ \ \ +H+(1+\frac{1}{H})\sum_{t=0}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t-1}-V^{*}_{h+1})(s,a)$
$\leq H+(1+\frac{1}{H})\sum_{t=0}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t-1}-V_{h+1}^{\pi^{t+1}})(s,a).$ (15)
For the fourth term, we can proceed in a similar way but using Lemma 15 to get
$\ \ \ \ \ \sum_{t=0}^{T-1}\frac{\chi_{h}^{t+1}(s,a)}{\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a)$
$\leq\sum_{k=t(s)}^{T-1}(\sum_{t=k}^{T-1}\frac{\chi_{h}^{t+1}(s,a)}{\tilde{n}_{h}^{t}(s,a)})\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a)$
$\leq 8log(T)\sum_{k=1}^{T-1}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a).$
Thus, we have
$\sum_{t=t(s)-1}^{T-1}\frac{5\chi_{h}^{t+1}(s,a)}{Hlog(T)\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a)$
$\leq\frac{40}{H}\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a).$ (16)
For the third term, combining Cauchy-Schwarz inequality, Lemmas 15 and 13, we get
$\ \ \ \ \ \sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)\sqrt{\frac{30}{\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{\tilde{n}_{h}^{t}(s,a)}}$
$\leq\sqrt{30\sum_{t=t(s)-1}^{T-1}\frac{\chi_{h}^{t+1}(s,a)}{\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)}\sqrt{\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)\frac{\zeta}{\tilde{n}_{h}^{t}(s,a)}}$
$\leq 44log(T)\sqrt{\zeta\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}.$ (17)
For the second term, we use Lemma 13 to get
$130H^{3}\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)\frac{log(T)\zeta}{\tilde{n}_{h}^{t}(s,a)}\leq 1040H^{3}log(T)^{2}\zeta.$ (18)
Finally, we combine eq. (15), eq. (16), eq. (17) and eq. (18) to conclude
$\hat{R}_{h}^{T}(s,a)\leq$ $44log(T)\sqrt{\zeta\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}+1041H^{3}log(T)^{2}\zeta$
$+(1+\frac{41}{H})\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a).$ (19)
Step 3: Replace $\chi_{h}^{t}$ with $\overline{p}_{h}^{t}$ in the upper-bound on $\hat{R}_{h}^{T}(s,a)$     Since $\mathcal{G}$ holds, we have
$\sqrt{\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}\leq\sqrt{2\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}+\sqrt{8\zeta}H$
and
$\ \ \ \ \ \sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a)$
$\leq(1+\frac{1}{H})\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a)+14H^{2}\zeta.$
Plugging the two inequalities in eq. (19) yields
$\hat{R}_{h}^{T}(s,a)\leq$ $63log(T)\sqrt{\zeta\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}+1754H^{3}log(T)^{2}\zeta$
$+(1+\frac{83}{H})\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a).$ (20)
Step 4: Upper-bound the step $h$ regret $\hat{R}_{h}^{T}$     We define
$\hat{R}_{h}^{T}\triangleq\sum_{s\in S}\sum_{t=0}^{T-1}\overline{p}_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s).$
We use the definition of $\mathring{\overline{V}}_{h}^{k}(s)$, Assumption 1 and the facts that $\mathcal{D}$ holds, $\mathring{\overline{Q}}_{h}^{k}\geq 0$ on $\mathcal{D}$ and that for all $x>1$ the following holds $\frac{1}{1-\frac{1}{4x}}\leq 1+\frac{1}{x}$ to get
$\ \ \ \ \ \sum_{t=0}^{T-1}\overline{p}_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)$
$=\sum_{t=0}^{(t(s)-2)\vee 0}\overline{p}_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)+\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)$
$\leq H(t(s)-1)+\frac{1}{1-\frac{1}{4H}}\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)+\frac{56}{3}H^{2}\zeta$
$\leq H\sqrt{T}+(1+\frac{1}{H})\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)+19H^{2}\zeta$
$\leq H\sqrt{T}+(1+\frac{1}{H})\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s)\pi_{h}^{t+1}(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s)+19H^{2}\zeta.$
Using Cauchy-Schwarz inequality,eq. (20) and the fact that the policy $\pi^{t}$ is deterministic, we obtain
$\hat{R}_{h}^{T}$ $\leq HS\sqrt{T}+(1+\frac{1}{H})\sum_{s\in S}\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s)\pi_{h}^{t+1}(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s)+19H^{2}S\zeta$
$=HS\sqrt{T}+(1+\frac{1}{H})\sum_{(s,a)\in S\times A}\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)+19H^{2}S\zeta$
$=HS\sqrt{T}+(1+\frac{1}{H})\sum_{(s,a)\in S\times A}\hat{R}_{h}^{T}(s,a)+19H^{2}S\zeta$
$\leq HS\sqrt{T}+126log(T)\sqrt{\zeta}\sum_{(s,a)\in S\times A}\sqrt{\sum_{t=t(s)-1}^{t-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}$
$\ \ \ \ \ +(1+\frac{167}{H})\sum_{(s,a)\in S\times A}\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a)$
$\ \ \ \ \ +3527SAH^{3}log(T)^{2}\zeta$
$\leq HS\sqrt{T}+3527SAH^{3}log(T)^{2}\zeta+(1+\frac{167}{H})\hat{R}_{h+1}^{T}$
$\ \ \ \ \ +126log(T)\sqrt{\zeta SA\sum_{(s,a)\in S\times A}\sum_{t=t(s)-1}^{t-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}.$ (21)
For the last inequality, we have
$\ \ \ \ \ \sum_{(s,a)\in S\times A}\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a)$
$=\sum_{(s,a,s^{\prime})\in S\times A\times S}\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(s^{\prime}|s,a)(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s^{\prime})$
$\leq\sum_{s^{\prime}\in S}\sum_{(s,a)\in S\times A}\sum_{t=0}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(s^{\prime}|s,a)(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s^{\prime})$
$=\sum_{s^{\prime}\in S}\sum_{t=0}^{T-1}\overline{p}_{h+1}^{t+1}(s^{\prime})(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s^{\prime})=\hat{R}_{h+1}^{T}.$
Step 5: Upper-bound the regret $\mathcal{R}^{T}$     Unfolding eq. (21), using Cauchy-Schwarz inequality, Lemma 12 and the fact that $\hat{R}^{T}_{H+1}=0$, we get
$\hat{R}_{1}^{T}$ $\leq\sum_{h=1}^{H}(1+\frac{167}{H})^{H-h}[126log(T)\sqrt{\zeta SA\sum_{(s,a)\in S\times A}\sum_{t=t(s)-1}^{t-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}$
$\ \ \ \ \ +HS\sqrt{T}+3527SAH^{3}log(T)^{2}\zeta]$
$\leq e^{167}126log(T)\sqrt{\zeta SAH\sum_{(s,a,h)\in S\times A\times[H]}\sum_{t=0}^{t-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}$
$\ \ \ \ \ +e^{167}H^{2}S\sqrt{T}+3527e^{167}SAH^{4}log(T)^{2}\zeta$
$=e^{167}126log(T)\sqrt{\zeta SAH\sum_{t=0}^{T-1}\mathbb{E}_{\pi^{t+1}}[(\sum_{h=1}^{H}r_{h}(s_{h},a_{h})-V_{1}^{\pi^{t+1}}(s_{1}))^{2}]}$
$\ \ \ \ \ +e^{167}H^{2}S\sqrt{T}+3527e^{167}SAH^{4}log(T)^{2}\zeta$
$\leq\tilde{\mathcal{O}}(\sqrt{\zeta H^{3}SAT})+\tilde{\mathcal{O}}(\zeta H^{4}SA)+\tilde{\mathcal{O}}(H^{2}S\sqrt{T}).$
Applying Lemma 9, we have
$V^{*}_{1}(s_{1})-V_{h}^{\pi^{t+1}}(s_{1})\leq\mathring{\overline{V}}^{t}_{1}(s_{1})-V_{h}^{\pi^{t+1}}(s_{1}).$
This allow us to conclude
$\mathcal{R}^{T}\leq\hat{R}_{1}^{T}\leq\tilde{\mathcal{O}}(\sqrt{\zeta H^{3}SAT})+\tilde{\mathcal{O}}(\zeta H^{4}SA)+\tilde{\mathcal{O}}(H^{2}S\sqrt{T}).$
Appendix K Preparation Lemmas
The proof of the following results can be seen in Ménard et al.[24]. For a deterministic policy $\pi$, we define Bellman-type equations for the variances as follows
$\sigma Q_{h}^{\pi}(s,a)\triangleq Var_{p_{h}}V_{h+1}^{\pi}(s,a)+p_{h}\sigma V_{h+1}^{\pi}(s,a)$
$\sigma V_{h}^{\pi}(s)\triangleq\sigma Q_{h}^{\pi}(s,\pi(s))$
$\sigma V_{H+1}^{\pi}(s)\triangleq 0,$
where $Var_{p_{h}}(f)(s,a)\triangleq\mathbb{E}_{s^{\prime}\sim p_{h}(\cdot|s,a)}[(f(s^{\prime})-p_{h}f(s,a))^{2}]$.
The definition above indicates that
$\sigma V_{1}^{\pi}(s_{1})=\sum_{h=1}^{H}\sum_{s,a}p_{h}^{\pi}(s,a)Var_{p_{h}}(V_{h+1}^{\pi})(s,a).$
We have the following lemmas.
Lemma 12.
For any deterministic policy $\pi$ and for all $h\in[H]$,
$\mathbb{E}_{\pi}[(\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})-Q_{h}(s_{h},a_{h}))^{2}|(s_{h},a_{h})=(s,a)]=\sigma Q_{h}^{\pi}(s,a).$
Particularly,
$\mathbb{E}_{\pi}[(\sum_{h=1}^{H}r_{h}(s_{h},a_{h})-V_{1}^{\pi}(s_{1}))^{2}]=\sigma V_{1}^{\pi}(s_{1})=\sum_{h=1}^{H}\sum_{s,a}p_{h}^{\pi}(s,a)Var_{p_{h}}(V_{h+1}^{\pi})(s,a).$
Lemma 13.
For $T\in\mathbb{N}^{*}$ and $(u_{t})_{t\in\mathbb{N}^{*}}$, for a sequence where $u_{t}\in[0,1]$ and $U_{t}\triangleq\sum_{l=1}^{t}u_{l}$, we get
$\sum_{t=1}^{T}\frac{u_{t+1}}{U_{t}\vee 1}\leq 4log(U_{T+1}+1).$
In particular if $T+1\geq 2$,
$\sum_{t=1}^{T}\frac{u_{t+1}}{U_{t}\vee 1}\leq 8log(T+1).$
Lemma 14.
For all $(s,a)\in S\times A$ the following holds
$\sum_{k=l}^{t}\chi_{h}^{k+1}(s,a)\tilde{\eta}_{h}^{k,l}(s,a)\leq(1+\frac{1}{H})\chi_{h}^{l}(s,a),$
$\sum_{k=t(s)}^{t}\tilde{\eta}_{h}^{t,k}(s,a)=1\ \ \ \ \ \text{if $n_{h}^{t}(s,a)>0$}.$
Lemma 15.
For all $(s,a)\in S\times A$ and $t\leq T-1$ (with $T\geq 2$), the following holds
$\chi_{h}^{l}(s,a)\sum_{k=l}^{t}\frac{\chi_{h}^{k+1}(s,a)}{\tilde{n}_{h}^{k}(s,a)}\leq\sum_{k=0}^{T-1}\frac{\chi_{h}^{k+1}(s,a)}{\tilde{n}_{h}^{k}(s,a)}=\sum_{k=t(s)-1}^{T-1}\frac{\chi_{h}^{k+1}(s,a)}{\tilde{n}_{h}^{k}(s,a)}\leq 8log(T).$
Lemma 16.
For $p,q\in\sum_{S}$, for $f,g:S\mapsto[0,b]$ two functions defined on $S$, we have that
$Var_{p}(f)\leq 2Var_{p}(g)+2bp|f-g|,\ \ \ Var_{p}(f^{2})\leq 2bVar_{p}(f),$
where we denote the absolute operator by $|f|(s)=|f(s)|$ for all $s\in S$.